### v1

# An Efficient Algorithm for Determining the Equivalence of Zero-one Reaction Networks

Abstract
Zero-one reaction networks play a crucial role in cell signaling. Determining the equivalence of reaction networks is a fundamental computational problem in the field of chemical reaction networks. In this work, we develop an efficient method for determining the equivalence of zero-one networks. The efficiency comes from several criteria for determining the equivalence of the steady-state ideals arising from zero-one networks, which helps for cutting down the expenses on computing Gröbner bases. Experiments show that our method can successfully classify over three million networks according to their equivalence in a reasonable time.
keywords:
Chemical reaction network, Gröbner basis, Steady-state ideal
1 Introduction
For the dynamical systems that arise from biochemical reaction networks, the following computational question is fundamental.
Question 1.
Which reaction networks have the same steady-state ideal?
The equivalence of steady-state ideals is one of the equivalence problem in the field of chemical reaction networks. The dynamical features of chemical reaction networks such as multistability, Hopf-bifurcation, and absolute concentration robustness (ACR) are linked to switch-like behavior, decision-making process, oscillation, and so on in cellular signaling [7, 23, 27, 9, 3, 25]. Usually, looking at the steady states is the first step for studying the dynamical behaviors of a chemical reaction system arising under mass-action kinetics. However, it is possible for two different networks to have the same steady-state ideal (notice that the steady states can be considered as real points located in the algebraic variety generated by the steady-state ideal). For instance, consider the two reaction networks:
$X_{1}+X_{2}\xrightarrow{\kappa_{1}}0,$ $X_{2}\xrightarrow{\kappa_{2}}X_{1}+X_{2},$ $0\xrightarrow{\kappa_{3}}X_{1}+X_{2},$ $0\xrightarrow{\kappa_{4}}X_{2};$ (1)
$X_{1}+X_{2}\xrightarrow{\kappa_{1}}X_{2},$ $X_{2}\xrightarrow{\kappa_{2}}X_{1},$ $0\xrightarrow{\kappa_{3}}X_{1},$ $0\xrightarrow{\kappa_{4}}X_{2}.$ (2)
They respectively have the following steady-state systems (see how to write down the system for a given network later in (8))
$\{-\kappa_{1}x_{1}x_{2}+\kappa_{2}x_{2}+\kappa_{3},-\kappa_{1}x_{1}x_{2}+\kappa_{3}+\kappa_{4}\};$ (3)
$\{-\kappa_{1}x_{1}x_{2}+\kappa_{2}x_{2}+\kappa_{3},-\kappa_{2}x_{2}+\kappa_{4}\}.$ (4)
It is straightforward to compute the reduced Gröbner bases [8] for the above two steady-state systems, and we find that in $\mathbb{Q}(\kappa)[x]$ they have the same reduced Gröbner basis
$\{\kappa_{2}x_{2}-\kappa_{4},\kappa_{1}\kappa_{4}x_{1}-\kappa_{2}\kappa_{3}-\kappa_{2}\kappa_{4}\}.$ (5)
In many cases, the above two networks are considered to be equivalent because when we only care about the dynamical behaviors determined by the steady states such as multistationarity, multistability, Hopf-bifurcation and ACR, we only need to study one of the networks that have the same steady-state ideal. Notice that relabeling the species or the reactions in a network naturally gives another kind of equivalence (see more details in [5] on the “dynamical equivalence” of networks). For instance, if we relabel the species $X_{1}$ and $X_{2}$ and we relabel the first and the second reactions in the network (1), then we get the following network
[…]
Of course, the network (6) should be considered to be equivalent to the network (1) and so, it is also equivalent to the network (2). Hence, in our context, we say two networks are equivalent if they have the same steady-state ideal after relabeling (see Definition 1).
Determining the equivalence is computationally challenging since a standard approach is to compute the reduced Gröbner bases. In a recent related work [12], the authors point out that there are few works on looking at the equivalence of the steady-state ideals, and we see that deriving explicit conditions for the equivalence is difficult in general. So, we might need to focus on a special class of networks for making this problem more feasible. The standard way (computing Gröbner bases [17, 15, 13, 14, 1, 16]) might be efficient for small networks such as one-species networks or two-reaction networks. However, in the studying of small networks, we usually need to deal with a substantial amount of networks. The idea of studying the small networks is motivated by the fact that many important dynamical features have inheritance from a large network to a smaller “subnetwork” such as multistability [20, 6], oscillation [2] and local bifurcations [4]. In many recent works, people are searching for the smallest networks from a big class of networks that admit these dynamical behaviors. In these studies, we need to first determine the equivalence of networks for making the searching process easier. For instance, Banaji and Boros [3] recently completely classify the smallest at-most-bimolecular networks that admit Hopf bifurcations (they contain three species and four reactions). Tang and Wang [27] find that the smallest zero-one networks that admit Hopf bifurcations are four-dimensional (they contain four species and five reactions). Kaihnsa, Nguyen and Shiu [21] prove that an at-most-bimolecular network admitting both multistationarity and ACR has at least three species and three reactions, and it is at least two-dimensional. The focus of this work is the zero-one networks since many biological significant networks in cell signaling are zero-one such as phosphorylation-dephosphorylation cycle [18], cell cycle [24], hybrid histidine kinase [26], and so on.
In this work, our main contribution is an efficient algorithm (Algorithm 1) for determining the equivalence of zero-one networks. The efficiency of the proposed new method comes from several criteria for determining the equivalence of the steady-state ideals arising from zero-one networks, which dramatically cuts down the expenses on computing Gröbner bases since many computations are avoided. We have successfully applied the new method to determining the equivalence of the smallest zero-one networks that admit multistationarity (these networks are three-dimensional, three-species and five-reaction [19]). The input contains over three millions such networks, and the new algorithm returns $32394$ networks after determining the equivalence. The computation is completed in a short time (see Table 1).
The rest of this paper is organized as follows. In Section 2, we review the basic concepts on reaction networks and the dynamical systems arising from mass-action kinetics. Also, we formally define the steady-state ideal and the equivalence of networks in our context. In Section 3, we prove a series of sufficient conditions (Theorems 13) for determining the equivalence of networks. In Section 4, we present the main algorithm (Algorithm 1) with two sub-algorithms (Algorithms 23) for determining the equivalence of zero-one networks. In Section 5, we illustrate how the main algorithm performs by applying it to the three-dimensional zero-one networks with three species and five reactions. Also, we explain the implementation details and present the computational timings.
2 Background
In this section, we briefly recall the standard notions and definitions for reaction networks and computational algebraic geometry, see [10] and [11] for more details.
A reaction network $G$ (or network for short) consists of a set of $s$ species $\{X_{1},X_{2},\dots,X_{s}\}$ and a set of $m$ reactions:
$\alpha_{1j}X_{1}+\dots+\alpha_{sj}X_{s}~{}\xrightarrow{\kappa_{j}}~{}\beta_{1j}X_{1}+\dots+\beta_{sj}X_{s},\;{\rm for}~{}j=1,2,\ldots,m,$ (7)
where all stoichiometric coefficients $\alpha_{ij}$ and $\beta_{ij}$ are non-negative integers, and we assume that $(\alpha_{1j},\ldots,\alpha_{sj})\neq(\beta_{1j},\ldots,\beta_{sj})$. Each $\kappa_{j}\in\mathbb{R}_{>0}$ is called a rate constant corresponding to the $j$-th reaction in (7). We say a reaction is a zero-one reaction, if the stoichiometric coefficients $\alpha_{ij}$ and $\beta_{ij}$ in (7) belong to $\{0,1\}$. We say a network (7) is a zero-one network if it only has zero-one reactions. For any reaction defined in (7), the stoichiometric vector is $(\beta_{1j}-\alpha_{1j},\dots,\beta_{sj}-\alpha_{sj})^{\top}\in\mathbb{Q}^{s}$. We say finitely many reactions are linearly independent if their corresponding stoichiometric vectors are linearly independent. We call the $s\times m$ matrix with $(i,j)$-entry equal to $\beta_{ij}-\alpha_{ij}$ the stoichiometric matrix of $G$, denoted by $\mathcal{N}$. The dimension of $G$ is defined as the rank of $\mathcal{N}$. We call the $s\times m$ matrix the reactant matrix of $G$ with $(i,j)$-entry equal to $\alpha_{ij}$, denoted by $\mathcal{Y}$.
Denote by $x_{1},\ldots,x_{s}$ the concentrations of the species $X_{1},\ldots,X_{s}$, respectively. Under the assumption of mass-action kinetics, we describe how these concentrations change in time by the following system of ODEs:
$\dot{x}\;=\;f(\kappa,x)\;:=\;\mathcal{N}v(\kappa,x)\;=\;\mathcal{N}\begin{pmatrix}\kappa_{1}\prod\limits_{i=1}^{s}x_{i}^{\alpha_{i1}}\\ \kappa_{2}\prod\limits_{i=1}^{s}x_{i}^{\alpha_{i2}}\\ \vdots\\ \kappa_{m}\prod\limits_{i=1}^{s}x_{i}^{\alpha_{im}}\\ \end{pmatrix},$ (8)
where $x=(x_{1},x_{2},\ldots,x_{s})^{\top}$, $v(\kappa,x)=(v_{1}(\kappa,x),\dots,v_{m}(\kappa,x))^{\top}$ and
$v_{j}(\kappa,x)\;:=\;\kappa_{j}\prod\limits_{i=1}^{s}x_{i}^{\alpha_{ij}}.$ (9)
By viewing the rate constants as a real vector of parameters $\kappa:=(\kappa_{1},\kappa_{2},\dots,\kappa_{m})^{\top}$, we have polynomials $f_{i}(\kappa,x)\in\mathbb{Q}(\kappa)[x]$, for $i\in\{1,\dots,s\}$. The polynomial ideal generated by $f$ in $\mathbb{Q}(\kappa)[x]$, denoted by $\langle f\rangle$, is called the steady-state ideal.
Definition 1.
Two reaction networks are equivalent if they have the same steady-state ideal in the sense of relabeling.
For any given rate-constant vector $\kappa^{*}\in{\mathbb{R}}^{m}_{>0}$, a steady state of (8) is a vector of concentrations $x^{*}\in\mathbb{R}_{\geq 0}^{s}$ such that $f(\kappa^{*},x^{*})=\vec{0}$, where $f(\kappa,x)$ is on the right-hand side of the ODEs (8), and $\vec{0}$ denotes the column vector whose coordinates are all zero. If all the coordinates of a steady state $x^{*}$ are strictly positive (i.e., $x^{*}\in\mathbb{R}_{>0}^{s}$), then we call $x^{*}$ a positive steady state. For the stoichiometric matrix $\mathcal{N}\in\mathbb{Q}^{s\times m}$, the positive flux cone of $\mathcal{N}$ is defined as
$\mathcal{F}^{+}(\mathcal{N})\;:=\;\{\gamma\in{\mathbb{R}}_{>0}^{m}|\mathcal{N}\gamma=\vec{0}\}.$ (10)
In applications, we usually only care about positive steady states of a network. It is well-known that a network admits positive steady states if and only if $\mathcal{F}^{+}(\mathcal{N})\neq\emptyset$.
3 Theorem
3.1 Sufficient conditions for vacuous ideals
For any matrix $\mathcal{A}$, we denote by $row_{i}(\mathcal{A})$ and $col_{i}(\mathcal{A})$ the $i$-th row and the $i$-th column of $\mathcal{A}$, respectively. For any vector $a$, we denote by $a_{i}$ the $i$-th coordinate of $a$.
Theorem 1.
For any network $G$ (7), let $f$ be the steady-state system defined as in (8). If there exists $a\in\mathbb{Q}^{s}$ such that
  • (1)
    $P:=\mathcal{N}^{\top}a\neq\vec{0}$,  and
  • (2)
    for any $i\in\{1,\dots,s\}$, if $P_{i}\neq 0$, then we have $col_{i}(\mathcal{Y})=\vec{0}$,
then $\langle f\rangle=\langle 1\rangle$ in ${\mathbb{Q}}(\kappa)[x]$.
Proof.
Let $v=(v_{1},\dots,v_{m})^{\top}$ be defined as in (9). By the condition (1), we have
$P^{\top}v=\sum_{i=1}^{s}a_{i}row_{i}(\mathcal{N})v.$ (11)
Note that $P\neq\vec{0}$. We denote by $P_{i_{1}},\dots,P_{i_{n}}$ all the non-zero coordinates in $P$. So, by (11), we have
$\sum_{i=1}^{s}a_{i}row_{i}(\mathcal{N})v=\sum_{j=1}^{n}P_{i_{j}}v_{i_{j}}.$ (12)
By the condition (2), for any $j\in\{1,\dots,n\}$, we have $col_{i_{j}}(\mathcal{Y})=\vec{0}$. Hence, by (9), for any $j\in\{1,\dots,n\}$,
$v_{i_{j}}=\kappa_{i_{j}}.$
Then, by (8) and by (12), we have
$\sum_{i=1}^{s}a_{i}f_{i}=\sum_{i=1}^{s}a_{i}row_{i}(\mathcal{N})v=\sum\limits_{j=1}^{n}P_{i_{j}}\kappa_{i_{j}}\in\langle f\rangle.$
Thus, we have $\langle f\rangle=\langle 1\rangle$ in ${\mathbb{Q}}(\kappa)[x]$. ∎
Example 1.
We illustrate how Theorem 2 works by the following network:
[…]
The corresponding reactant matrix is
$\mathcal{Y}=\begin{pmatrix}1&0&0\\ 0&1&0\end{pmatrix}.$ (13)
The corresponding stoichiometric matrix is
$\mathcal{N}=\begin{pmatrix}-1&1&0\\ 1&-1&1\end{pmatrix}.$
Let $a=(1,1)^{\top}$. Then, we have $P:=\mathcal{N}^{\top}a=(0,0,1)^{\top}$. Note that $P_{3}\neq 0$. By (13), $col_{3}(\mathcal{Y})=\vec{0}$. Hence, by Theorem 1, we have $\langle f\rangle=\langle 1\rangle$.
A row (column) of a matrix is called a non-zero row (column) if there exists a non-zero element in this row (column). If a row (column) has both positive and negative elements, we say the row (column) changes signs.
Theorem 2.
For any $d$-dimensional zero-one network $G$ (7), let $f$ be the steady-state system defined as in (8). We denote by $c$ the number of zero columns of the reactant matrix $\mathcal{Y}$. Assume that all the rows of the stoichiometric matrix $\mathcal{N}$ are non-zero and change signs. Then, we have the following statements.
  1. 1.
    There is at least one non-zero column in $\mathcal{Y}$ (i.e., $c<m$).
  2. 2.
    If $c=m-1$ and $d\geq 2$, then $\langle f\rangle=\langle 1\rangle$ in ${\mathbb{Q}}(\kappa)[x]$.
  3. 3.
    If $c=m-2$ and $d=s\geq 3$, then $\langle f\rangle=\langle 1\rangle$ in ${\mathbb{Q}}(\kappa)[x]$.
  4. 4.
    If $c=0$, then $\langle f\rangle\neq\langle 1\rangle$ in ${\mathbb{Q}}(\kappa)[x]$.
Proof.
  1. 1.
    Recall that any stoichiometric coefficients $\alpha_{ij}$ and $\beta_{ij}$ are defined as in (7). By the definition of zero-one network, for any $(i,j)\in\{1,\dots,s\}\times\{1,\dots,m\}$, we have $\beta_{ij}\in\{0,1\}$. So, if there exists $(i,j)$ such that $\alpha_{ij}=0$, then the $(i,j)$-entry of $\mathcal{N}$ (i.e., $\beta_{ij}-\alpha_{ij}$) can only be 00 or $1$. Hence, if all entries of $\mathcal{Y}$ are 00 (i.e., $c=m$), then all entries of $\mathcal{N}$ are 00 or $1$. It is contrary to the assumption that all the rows of $\mathcal{N}$ are non-zero and change signs. Therefore, there is at least one non-zero column in $\mathcal{Y}$ (i.e., $c<m$).
  2. 2.
    Since all the rows of $\mathcal{N}$ are non-zero and change signs, there exists the element $-1$ in every row of $\mathcal{N}$. By the definition of zero-one network, we have $\alpha_{ij}\in\{0,1\}$ and $\beta_{ij}\in\{0,1\}$. Hence, if there exists an $(i,j)$-entry of $\mathcal{N}$ such that the entry is $-1$ (i.e., $\beta_{ij}-\alpha_{ij}=-1$), then $\alpha_{ij}=1$. Hence, there exists the element $1$ in every row of $\mathcal{Y}$. Since the number of zero columns is $m-1$, we can assume that the first $m-1$ columns of $\mathcal{Y}$ are the zero vector. Note that there exists the element $1$ in every row of $\mathcal{Y}$. So, all the coordinates of the the $m$-th column of $\mathcal{Y}$ must be $1$. Then, we have
    $\mathcal{Y}=\begin{pmatrix}0&\cdots&0&1\\ &\ddots&&\vdots\\ 0&\cdots&0&1\end{pmatrix}.$
    By the definition of zero-one network, if there exists $(i,j)$ such that $\alpha_{ij}=0$, then the $(i,j)$-entry of $\mathcal{N}$ (i.e., $\beta_{ij}-\alpha_{ij}$) can only be 00 or $1$. Hence, every coordinate of the first $m-1$ columns of $\mathcal{N}$ can only be 00 or $1$. Note that there exists $-1$ in every row vector of $\mathcal{N}$. So, all the coordinates of the the $m$-th column of $\mathcal{N}$ are $-1$. Thus, $\mathcal{N}$ has the following form
    $\mathcal{N}=\begin{pmatrix}&&&-1\\ &\mathcal{A}_{1}&&\vdots\\ &&&-1\end{pmatrix},$ (14)
    where all the entries of the $s\times(m-1)$ submatrix $\mathcal{A}_{1}$ are 00 or $1$. Recall that for any $i\in\{1,\dots,s\}$, we denote by $row_{i}(\mathcal{N})$ the $i$-th row of $\mathcal{N}$. Since $d\geq 2$ (note here, ${\tt rank}({\mathcal{N}})=d$), there exist $i,j\in\{1,\dots,s\}$ such that $P^{\top}:=row_{i}(\mathcal{N})-row_{j}(\mathcal{N})\neq\vec{0}^{\top}$. So, if $P_{k}\neq 0$, then by (14), we have $k\in\{1,\dots,m-1\}$. Note that for any $k\in\{1,\dots,m-1\}$, the $k$-th column vector of $\mathcal{Y}$ is a zero vector. Therefore, by Theorem 1, we have $\langle f\rangle=\langle 1\rangle$.
  3. 3.
    Assume that the first $m-2$ columns of $\mathcal{Y}$ are the zero vector. Then, by the definition of zero-one network, we have
    $\mathcal{Y}=\begin{pmatrix}0&\cdots&0&\\ &\ddots&&\mathcal{A}_{2}\\ 0&\cdots&0&\end{pmatrix},$
    where the entries of the $s\times 2$ submatrix $\mathcal{A}_{2}$ are 00 or $1$. Then, similar to the proof of (2), by the definition of zero-one network, we have
    $\mathcal{N}=\begin{pmatrix}\mathcal{A}_{3}&\mathcal{A}_{4}\end{pmatrix},$ (15)
    where the entries of the $s\times(m-2)$ submatrix $\mathcal{A}_{3}$ are 00 or $1$. Since all rows of $\mathcal{N}$ are non-zero and change signs, there exists the element $-1$ in every row of $\mathcal{N}$. Hence, the rows of the $s\times 2$ submatrix $\mathcal{A}_{4}$ must be in the set
    $\{(0,-1),(-1,0),(1,-1),(-1,1),(-1,-1)\}.$ (16)
    • (i)
      Assume that there exist $k<s$ different row vectors in (16) appearing in $\mathcal{A}_{4}$. Since $k<s$, at least two rows in $\mathcal{A}_{4}$ are the same. Assume that the $i$-th and the $j$-th ($i,j\in\{1,\dots,s\}$) rows in $\mathcal{A}_{4}$ are the same. Define $P^{\top}:=row_{i}(\mathcal{N})-row_{j}(\mathcal{N})$. Then, we have $P_{m-1}=P_{m}=0$. Since ${\tt rank}(\mathcal{N})=d=s$, we have $P\neq\vec{0}$. Therefore, for any $k\in\{1,\dots,m\}$, if $P_{k}\neq 0$, we have $k\in\{1,\dots,m-2\}$. Note that for any $k\in\{1,\dots,m-2\}$, the $k$-th column of $\mathcal{Y}$ is the zero vector. Hence, by Theorem 1, we have $\langle f\rangle=\langle 1\rangle$.
    • (ii)
      Assume that there exist $s$ different row vectors in (16) appearing in $\mathcal{A}_{4}$. Note that any matrix consisting of $s\geq 3$ different row vectors in (16) is rank-two. Hence, ${\rm rank}(\mathcal{A}_{4})=2$. Note that there are $s\geq 3$ rows in $\mathcal{A}_{4}$. So, there exists $a\in\mathbb{Q}^{s}\setminus\{\vec{0}\}$ such that $\mathcal{A}_{4}^{\top}a=\vec{0}$. Recall that $a=(a_{1},\dots,a_{s})^{\top}$. By (15), $\mathcal{A}_{4}^{\top}a$ is equal to the last two coordinates of $P^{\top}:=\sum\limits_{i=1}^{s}a_{i}row_{i}(\mathcal{N})$. Hence, we have $P_{m-1}=P_{m}=0$. Note that ${\rm rank}(\mathcal{N})=d=s$. So, $P\neq\vec{0}$ since $a\neq\vec{0}$. Hence, for any $k\in\{1,\dots,m\}$, if $P_{k}\neq 0$, then we have $k\in\{1,\dots,m-2\}$. Note that for any $k\in\{1,\dots,m-2\}$, the $k$-th column of $\mathcal{Y}$ is the zero vector. Therefore, by Theorem 1, we have $\langle f\rangle=\langle 1\rangle$.
  4. 4.
    Since there is no zero columns in $\mathcal{Y}$, by (9), for any $i\in\{1,\dots,m\}$, $v_{i}=\kappa_{i}\prod_{j=1}^{s}x^{\alpha_{ji}}$, where $\sum\limits_{j=1}^{s}\alpha_{ji}\geq 0$ since $\alpha_{ji}\in\{0,1\}$. Hence, by (8), all terms in $f$ contain the variables $x_{i}$’s. So, in ${\mathbb{Q}}(\kappa)[x]$, $1\notin\langle f\rangle$. Hence, we have $\langle f\rangle\neq\langle 1\rangle$.
Example 2.
We illustrate how Theorem 2 works by the following examples.
(1) Consider the following zero-one network:
[…]
$0\displaystyle 0\xrightarrow{\kappa_{4}}X_{1},\;\;X_{1}+X_{2}+X_{3}\xrightarrow{\kappa_{5}}$ 0 .
The stoichiometric matrix $\mathcal{N}$ is
$\begin{pmatrix}0&0&0&1&-1\\ 0&1&1&0&-1\\ 1&0&1&0&-1\end{pmatrix}.$
Note that ${\rm rank}(\mathcal{N})=3$. The reactant matrix $\mathcal{Y}$ is
$\begin{pmatrix}0&0&0&0&1\\ 0&0&0&0&1\\ 0&0&0&0&1\end{pmatrix}.$
Note that there are $4$ zero columns in $\mathcal{Y}$ (i.e., $c=4=m-1$). Hence, by Theorem 2 (2), we have $\langle f\rangle=\langle 1\rangle$. Note that
$f_{1}=$ $\kappa_{4}-\kappa_{5}x_{1}x_{2}x_{3},$
$f_{2}=$ $\kappa_{2}+\kappa_{3}-\kappa_{5}x_{1}x_{2}x_{3},$
$f_{3}=$ $\kappa_{1}+\kappa_{3}-\kappa_{5}x_{1}x_{2}x_{3}.$
It is straightforward to check that $\frac{f_{1}-f_{2}}{\kappa_{4}-\kappa_{2}-\kappa_{3}}=1$.
(2) Consider the following zero-one network:
[…]
$X_{2}\xrightarrow{\kappa_{4}}X_{1},\;\;X_{1}+X_{2}+X_{3}\xrightarrow{\kappa_{5}}0.$
The stoichiometric matrix $\mathcal{N}$ is
$\begin{pmatrix}0&0&0&1&-1\\ 0&1&1&-1&-1\\ 1&0&1&0&-1\end{pmatrix}.$
Note that ${\rm rank}(\mathcal{N})=3$, i.e., $d=s=3$. The reactant matrix $\mathcal{Y}$ is
$\begin{pmatrix}0&0&0&0&1\\ 0&0&0&1&1\\ 0&0&0&0&1\end{pmatrix}.$
Note that there are $3$ zero columns in $\mathcal{Y}$ (i.e., $c=3=m-2$). Hence, by Theorem 2 (3), we have $\langle f\rangle=\langle 1\rangle$. Note that
$f_{1}=$ $\kappa_{4}x_{2}-\kappa_{5}x_{1}x_{2}x_{3},$
$f_{2}=$ $\kappa_{2}+\kappa_{3}-\kappa_{4}x_{2}-\kappa_{5}x_{1}x_{2}x_{3},$
$f_{3}=$ $\kappa_{1}+\kappa_{3}-\kappa_{5}x_{1}x_{2}x_{3}.$
It is straightforward to check that $\frac{f_{1}+f_{2}-2f_{3}}{-2\kappa_{1}+\kappa_{2}-\kappa_{3}}=1$.
(3) Consider the following zero-one network:
$X_{1}\xrightarrow{\kappa_{1}}X_{3},\;\;X_{1}\xrightarrow{\kappa_{2}}X_{2},\;\;X_{1}\xrightarrow{\kappa_{3}}X_{2}+X_{3},$
$X_{2}\xrightarrow{\kappa_{4}}X_{1},\;\;X_{1}+X_{2}+X_{3}\xrightarrow{\kappa_{5}}0.$
The stoichiometric matrix $\mathcal{N}$ is
$\begin{pmatrix}-1&-1&-1&1&-1\\ 0&1&1&-1&-1\\ 1&0&1&0&-1\end{pmatrix}.$
The reactant matrix $\mathcal{Y}$ is
$\begin{pmatrix}1&1&1&0&1\\ 0&0&0&1&1\\ 0&0&0&0&1\end{pmatrix}.$
Note that there is no zero columns in $\mathcal{Y}$ (i.e., $c=0$). Hence, by Theorem 2 (4), we have $\langle f\rangle\neq\langle 1\rangle$. Note that
$f_{1}=$ $-\kappa_{1}x_{1}-\kappa_{2}x_{1}-\kappa_{3}x_{1}+\kappa_{4}x_{2}-\kappa_{5}x_{1}x_{2}x_{3},$
$f_{2}=$ $\kappa_{2}x_{1}+\kappa_{3}x_{1}-\kappa_{4}x_{2}-\kappa_{5}x_{1}x_{2}x_{3},$
$f_{3}=$ $\kappa_{1}x_{1}+\kappa_{3}x_{1}-\kappa_{5}x_{1}x_{2}x_{3}.$
It is straightforward to check that $1\notin\langle f\rangle$.
3.2 Sufficient conditions for equivalence
Theorem 3.
For any two networks $G$ and $G^{^{\prime}}$, let $\mathcal{N}$ and $\mathcal{N}^{^{\prime}}$ be the corresponding stoichiometric matrices, and let $\mathcal{Y}$ and $\mathcal{Y}^{^{\prime}}$ be the reactant matrices. Let $f$ and $f^{^{\prime}}$ be the steady-state systems defined as in (8). If
  • (1)
    $\mathcal{Y}=\mathcal{Y}^{^{\prime}}$, and
  • (2)
    for any $i\in\{1,\dots,s\}$, we have
    $row_{i}(\mathcal{N})\in\operatorname{span}_{\mathbb{Q}}\{row_{1}(\mathcal{N}^{^{\prime}}),\dots,row_{s}(\mathcal{N}^{^{\prime}})\},$
    and
    $row_{i}(\mathcal{N}^{^{\prime}})\in\operatorname{span}_{\mathbb{Q}}\{row_{1}(\mathcal{N}),\dots,row_{s}(\mathcal{N})\},$
where $\operatorname{span}_{\mathbb{Q}}\{\cdot\}$ means the rational vector space spanned by a set of rational vectors, then $\langle f\rangle=\langle f^{^{\prime}}\rangle$ in ${\mathbb{Q}}(\kappa)[x]$.
Proof.
Let $v=(v_{1},\dots,v_{m})^{\top}$ and $v^{^{\prime}}=(v_{1}^{^{\prime}},\dots,v_{m}^{^{\prime}})^{\top}$ be defined as in (9) corresponding to $G$ and $G^{\prime}$ respectively, where
$v_{i}=\kappa_{i}\prod_{j=1}^{s}x_{j}^{\alpha_{ji}},\;\;v_{i}^{^{\prime}}=\kappa_{i}\prod_{j=1}^{s}x_{j}^{\alpha_{ji}^{^{\prime}}}.$ (17)
By $\mathcal{Y}=\mathcal{Y}^{^{\prime}}$, for any $i\in\{1,\dots,s\}$ and for any $j\in\{1,\dots,s\}$, $\alpha_{ij}=\alpha_{ij}^{^{\prime}}$. Hence, by (17), we have $v=v^{^{\prime}}$. So, the conclusion follows from (8) and the condition (2). ∎
Example 3.
Consider the following two networks $G$ and $G^{^{\prime}}$.
$G:$ $\quad X_{1}+X_{2}\xrightarrow{\kappa_{1}}X_{2},$ $X_{2}\xrightarrow{\kappa_{2}}0,$ $0\xrightarrow{\kappa_{3}}X_{1}+X_{2},$
$G^{^{\prime}}:$ $\quad X_{1}+X_{2}\xrightarrow{\kappa_{1}}X_{2},$ $X_{2}\xrightarrow{\kappa_{2}}X_{1},$ $0\xrightarrow{\kappa_{3}}X_{2}.$
Note that the two networks $G$ and $G^{^{\prime}}$ have the same set of reactants. Hence, the reactant matrices $\mathcal{Y}$ and $\mathcal{Y}^{^{\prime}}$ are equal. The stoichiometric matrices $\mathcal{N}$ and $\mathcal{N}^{^{\prime}}$ are as follows.
$\mathcal{N}=\begin{pmatrix}-1&0&1\\ 0&-1&1\end{pmatrix},\quad\mathcal{N}^{^{\prime}}=\begin{pmatrix}-1&1&0\\ 0&-1&1\end{pmatrix}.$
Note that $row_{1}(\mathcal{N})=row_{1}(\mathcal{N}^{^{\prime}})+row_{2}(\mathcal{N}^{^{\prime}})$, $row_{2}(\mathcal{N})=row_{2}(\mathcal{N}^{^{\prime}})$ and $row_{1}(\mathcal{N}^{^{\prime}})=row_{1}(\mathcal{N})-row_{2}(\mathcal{N})$. So, by Theorem 3, the ideals generated by $f$ and $f^{^{\prime}}$ are equal.
Remark 1.
We remark that although Theorem 1 and Theorem 3 are straightforward to prove, the hypotheses in these theorems can not be directly checked by a computer, while Theorem 2 is explicit enough. In the next section, we will solve this issue by Algorithms 2 and 3.
4 Algorithm
In this section, we propose a main algorithm (Algorithm 1) to determine the equivalence of zero-one networks with two sub-algorithms (Algorithm 2 and 3). The goal of the main algorithm is to efficiently classify all $d$-dimensional $s$-species $m$-reaction zero-one networks according to the equivalence determined by their steady-state ideals. The correctness of Algorithm 1 follows from Theorem 3 and the correctness of the two sub-algorithms.
Algorithm 1.
DetermineEquivalence
Input.
The number of the species $s$, the number of the reactions $m$ and the dimension $d$
Output.
Finitely many classes of $d$-dimensional $s$-species $m$-reaction zero-one networks such that the networks in the same class have the same steady-state ideal
Step 1.
Enumerate all $d$-dimensional $s$-species $m$-reaction zero-one networks.
Step 2.
Pick the networks satisfying the following two conditions simultaneously.
  1. 1.
    All the rows of the stoichiometric matrix $N$ are non-zero and change signs.
  2. 2.
    The set $\mathcal{F}^{+}(\mathcal{N})$ defined as in (10) is not empty.
(##\##By this step, we remove the networks definitely admitting no positive steady states.)
Step 3.
For each remaining network, do the following steps:
Step 3.1.
If $c=m-1$ and $d\geq 2$, then delete the network.
Step 3.2.
If $c=m-2$ and $d=s\geq 3$, then delete the network.
(##\##Note here, $c$ denotes the number of zero columns of the reactant matrix $\mathcal{Y}$.)
Step 4.
We call CoefficientsForVacuous($s$, $d$), and we get a set $S\subset\mathbb{Q}^{s}$.
Step 5.
For each remaining network, if there exists $a\in S$ such that
  • (1)
    $\mathcal{N}^{\top}a\neq 0$, and
  • (2)
    for any $i\in\{1,\dots,s\}$, if the $i$-th coordinate of $\mathcal{N}^{\top}a$ is non-zero, then $col_{i}(\mathcal{Y})=\vec{0}$,
then delete the network.
Step 6.
Delete the same networks after relabeling the species.
Step 7.
We call CoefficientsForEquivalence($s$, $d$), and we get a set $S_{a}\subset\mathbb{Q}^{s}$.
Step 8.
For each pair of networks $G$ and $G^{^{\prime}}$ from the remaining networks in Step 6, do the following steps.
Step 8.1.
If there exists a permutation on the rows of $\mathcal{Y}$ and $\mathcal{Y}^{^{\prime}}$ such that $\mathcal{Y}$ and $\mathcal{Y}^{^{\prime}}$ are equal after permutating their rows in the same way, then permutate the rows of $\mathcal{N}$ and $\mathcal{N}^{^{\prime}}$ in the same way, and go to Step 8.2. Otherwise, consider the next pair of networks.
Step 8.2.
Check if $\mathcal{N}$ and $\mathcal{N}^{^{\prime}}$ satisfy the following condition: for any $i\in\{1,\dots,s\}$, there exist $(a_{i1}\dots,a_{is})^{\top}\in S_{a}$ and $(b_{i1},\dots,b_{is})^{\top}\in S_{a}$ such that
$\qquad\quad\enspace row_{i}(\mathcal{N}^{^{\prime}})=\sum\limits_{j=1}^{s}a_{ij}row_{j}(\mathcal{N}),\;row_{i}(\mathcal{N})=\sum\limits_{j=1}^{s}b_{ij}row_{j}(\mathcal{N}^{^{\prime}}).$
If yes, then put the networks $G$ and $G^{^{\prime}}$ into the same class. If not, then consider the next pair of networks.
Step 9.
For every class generated in Step 8, choose the first network as the representative element. Calculate the reduced Gröbner bases for these representative networks. Delete the networks whose reduced Gröbner base are $\{1\}$ or have a monomial as an element.
Step 10.
Combine the networks with the same reduced Gröbner basis into the same class. Choose the first network as the representative element. Return these representative elements.
Algorithm 2.
CoefficientsForVacuous
Input.
The number of species $s$ and the dimension $d$
Output.
A set $S\subset\mathbb{Q}^{s}$ such that for any $d$-dimensional $s$-species zero-one network, if there exists $a\in\mathbb{Q}^{s}$ such that the conditions (1) and (2) of Theorem 1 hold, then there exists $\tilde{a}\in S$ satisfying the conditions (1) and (2) of Theorem 1
Step 1.
Enumerate all $s$-species zero-one reactions.
Step 2.
Let $S:=\emptyset$. For every $n\in\{1,\dots,d-1\}$, choose $n$ linearly independent reactions from the reactions enumerated in Step 1. For every choice, do the following steps.
Step 2.1.
Denote by
$\mathcal{M}=\begin{pmatrix}M_{1}\\ \vdots\\ M_{n}\end{pmatrix}$
an $n\times s$ matrix, where each row $M_{j}$ is the transpose of the stoichiometric vector of a reaction from the chosen $n$ independent reactions.
Step 2.2.
Calculate the basis of the null space of $\mathcal{M}$ and add all the vectors in the basis into the set $S$.
(##\##Note here, all these vectors in the basis are in $\mathbb{Q}^{s}$.)
Step 3.
Return the set $S$.
Proof of the Correctness of Algorithm 2. Define
$\Delta_{1}$ :=$:=\{i\in\{1,\dots,m\}|col_{i}(\mathcal{Y})\neq\vec{0}\},$ (18)
$\Delta_{2}$ :=$:=\{1,\dots,m\}\setminus\Delta_{1}.$ (19)
Let $M$ and $M_{k}$ ($k\in\{1,2\}$) be the real vector spaces spanned by the columns of ${\mathcal{N}}$ and the set of column vectors $\{col_{i}({\mathcal{N}})|i\in\Delta_{k}\}$, respectively. By (18) and the condition (2) of Theorem 1, we have
$a\in M_{1}^{\perp}.$
Since $\mathcal{N}^{\top}a\neq\vec{0}$ (the condition (1) of Theorem 1), we have $a\not\in M^{\perp}$. Since $M_{1}\subset M$, we have ${\tt dim}(M_{1})<{\tt dim}(M)=d$. So, by Step 2 of Algorithm 2, there exists $B\subset S$ such that $B$ is the basis of $\mathcal{M}^{\perp}_{1}$. Notice that there exists $\tilde{a}\in B$ such that $\tilde{a}\not\in M^{\perp}$, i.e., $\mathcal{N}^{\top}\tilde{a}\neq\vec{0}$. Otherwise, we have $B\subset M^{\perp}$ and hence, we have $M_{1}^{\perp}\subset M^{\perp}$, which is contrary to the fact that $a\in M_{1}^{\perp}\backslash M^{\perp}$. So, the conditions (1) and (2) of Theorem 1 hold for $\tilde{a}\in{\mathbb{Q}}^{s}$.
Example 4.
We illustrate how Algorithm 2 works for $s=d=3$.
  1. Step 1.
    Enumerate all three-species zero-one reactions (there are $56$ reactions in total).
  2. Step 2.
    Let $S:=\emptyset$. Note that $d=3$. For every $n\in\{1,2\}$, choose $n$ linearly independent reactions from the $56$ reactions enumerated in Step 1 and for every choice, do the following steps (below we only give details on the first round).
    1. Step 2.1
      Suppose $(0,0,1)$ is the stoichiometric vector corresponding to the first reaction from the $56$ reactions. Let $\mathcal{M}:=(0,0,1)$.
    2. Step 2.2
      The basis of the null space of $\mathcal{M}$ is $\{(1,0,0)^{\top},(0,1,0)^{\top}\}$. Add the vectors in the basis into the set $S$.
    Similarly, we do Steps 2.1 and 2.2 for all the other choices of reactions. Then, we get the set $S$:
    $\{0,1\}^{3}\cup(\{-1,1\}^{s}\times\{2\})\cup$
    ($(\{-1,1\}\times\{-2,2\}\times\{1\})\cup(\{-2,2\}\times\{-1,1\}\times\{1\})\cup$
    $\{(-1,-1,1)^{\top},(1,-1,1)^{\top},(-1,1,1)^{\top},(-1,0,1)^{\top},$
    $(-1,1,0)^{\top},(0,-1,1)^{\top}\}\setminus(0,0,0)^{\top}.$ (20)
Algorithm 3.
CoefficientsForEquivalence
Input.
The number of species $s$ and the dimension $d$
Output.
A set $S_{a}\subset\mathbb{Q}^{s}$ such that if two $d$-dimensional $s$-species zero-one networks satisfy
  • (1)
    ${\mathcal{Y}}={\mathcal{Y}}^{^{\prime}}$, and
  • (2)
    for any $i\in\{1,\dots,s\}$,
    $row_{i}(\mathcal{N}^{^{\prime}})$ $\in\operatorname{span}_{\mathbb{Q}}\{row_{1}(\mathcal{N}),\dots,row_{s}(\mathcal{N})\},$ (21)
then for any $i\in\{1,\dots,s\}$, there exists $(a_{i1},\dots,a_{is})^{\top}\in S_{a}$ such that
$row_{i}(\mathcal{N}^{^{\prime}})=\sum\limits_{j=1}^{s}a_{ij}row_{j}(\mathcal{N})$
Step 1.
Enumerate all possible zero-one reactions with $s$ species (here, we only enumerate one direction in a pair of reversible reactions). Write the corresponding stoichiometric vectors into a set $E\subset\mathbb{Q}^{s}$ (here, it is possible for different reactions to have the same stoichiometric vector, and we only put different vectors into the set $E$).
Step 2.
Let $S_{a}:=\emptyset$. For any $d$ column vectors from $E$, do the following steps.
Step 2.1.
Let $\mathcal{A}$ be the $s\times d$ matrix consisting of the $d$ column vectors. If $\mathcal{A}$ has any zero rows, then stop and go to the next group of $d$ vectors chosen from $E$.
Step 2.2.
Denote by $\vec{1}$ and $-\vec{1}$ the column vectors whose coordinates are all $1$ and $-1$, respectively. Define
$\Lambda_{1}$ :=$:=\{i\in\{1,\dots,d\}|col_{i}(\mathcal{A})=\vec{1}\},$ (22)
$\Lambda_{2}$ :=$:=\{i\in\{1,\dots,d\}|col_{i}(\mathcal{A})=-\vec{1}\},$ (23)
$\Lambda_{3}$ :=$:=\{1,\dots,d\}\setminus(\Lambda_{1}\cup\Lambda_{2}).$ (24)
Define a set of vectors
$F:=\{0,1\}^{|\Lambda_{1}|}\times\{-1,0\}^{|\Lambda_{2}|}\times\{-1,0,1\}^{|\Lambda_{3}|}\subset\mathbb{Q}^{d}.$ (25)
(##\##Notice here $|\Lambda_{1}|+|\Lambda_{2}|+|\Lambda_{3}|=d$.)
For any column vector $b$ in the set $F$, do the following steps.
Step 2.2.1.
Solve $a\in\mathbb{Q}^{s}$ from $\mathcal{A}^{\top}a=b$.
Step 2.2.2.
If there exists finitely many solutions, then add the solutions into the set $S_{a}$. If there exists infinitely many solutions, then add one of the solutions into the set $S_{a}$. If there is no solutions, then go to the next vector $b\in F$.
Step 3.
Return the set $S_{a}$.
Proof of the Correctness of Algorithm 3. Consider two $d$-dimensional $s$-species zero-one networks $G$ and $G^{^{\prime}}$. Assume that their reactant matrices are equal, i.e., $\mathcal{Y}=\mathcal{Y}^{^{\prime}}$. First, we prove a claim.
Claim. If there exists $\mathcal{A}$ in Step 2.1 of Algorithm 3 such that $\mathcal{A}=\mathcal{N}$, then for any $j\in\{1,\dots,s\}$, there exists $b\in F$ (defined as in (25)) such that $row_{j}(\mathcal{N}^{^{\prime}})^{\top}=b$.
In fact, by (22), for any $i\in\Lambda_{1}$, all the coordinates of $col_{i}(\mathcal{A})$ are $1$. Then, since $\mathcal{N}=\mathcal{A}$, for any $i\in\Lambda_{1}$, all the coordinates of $col_{i}(\mathcal{N})$ are $1$. By the definition of zero-one network, all the coordinates of $col_{i}(\mathcal{Y})$ are 00. Then, since $\mathcal{Y}=\mathcal{Y}^{^{\prime}}$, all the coordinates of $col_{i}(\mathcal{Y}^{^{\prime}})$ are 00. Thus, by the definition of zero-one network again, we have all the coordinates of $col_{i}(\mathcal{N}^{^{\prime}})$ are 00 or $1$. Hence, for any $j\in\{1,\dots,s\}$ and for any $i\in\Lambda_{1}$, the $i$-th coordinate of $row_{j}(\mathcal{N}^{^{\prime}})$ is 00 or $1$. Similarly, we can derive that for any $j\in\{1,\dots,s\}$ and for any $i\in\Lambda_{2}$, the $i$-th coordinate of $row_{j}(\mathcal{N}^{^{\prime}})$ is $-1$ or 00. Note that by the definition of zero-one network, for any $j\in\{1,\dots,s\}$ and for any $i\in\Lambda_{3}$, the $i$-th coordinate of $row_{j}(\mathcal{N}^{^{\prime}})$ is $-1$, 00 or $1$. Hence, by (25), for any $j\in\{1,\dots,s\}$, $row_{j}(\mathcal{N}^{^{\prime}})^{\top}\in F$. So, for any $j\in\{1,\dots,s\}$, there exists $b\in F$ such that $row_{j}(\mathcal{N}^{^{\prime}})^{\top}=b$.
Below, we prove the correctness of Algorithm 3 by discussing two cases.
  1. Case 1.
    Assume that all the columns of $\mathcal{N}$ are in the set $E$ generated in Step 1 of Algorithm 3.
    1. (i)
      Assume that all the columns of $\mathcal{N}$ are different. Since all the columns in $E$ are different. by the choice of $\mathcal{A}$ in Step 2 of Algorithm 3, there exists $\mathcal{A}$ such that $\mathcal{A}=\mathcal{N}$. By (21), for any $i\in\{1,\dots,s\}$, there exists $a^{^{\prime}}:=(a_{i1}^{^{\prime}},\dots,a_{is}^{^{\prime}})^{\top}\in\mathbb{Q}^{s}$ such that
      $row_{i}(\mathcal{N}^{^{\prime}})=\sum\limits_{j=1}^{s}a_{ij}^{^{\prime}}row_{j}(\mathcal{N}).$ (26)
      Hence, by (26), for any $i\in\{1,\dots,s\}$, we have
      $\mathcal{N}^{\top}a^{^{\prime}}=row_{i}(\mathcal{N}^{^{\prime}})^{\top}.$ (27)
      By $\mathcal{A}=\mathcal{N}$ and by the claim, for any $i\in\{1,\dots,s\}$, there exists $b\in F$ such that $row_{i}(\mathcal{N}^{^{\prime}})^{\top}=b$. Hence, by $\mathcal{N}=\mathcal{A}$ and (27), we have $\mathcal{A}^{\top}a^{^{\prime}}=b$, i.e., there exists a solution to the equation $\mathcal{A}^{\top}a=b$. By Step 2.2.2, there exists $a\in S_{a}$ such that $\mathcal{A}^{\top}a=b$. Hence, for any $i\in\{1,\dots,s\}$, by $\mathcal{A}=\mathcal{N}$ and $row_{i}(\mathcal{N}^{^{\prime}})^{\top}=b$, we have $\mathcal{N}^{\top}a=row_{i}(\mathcal{N}^{^{\prime}})^{\top}$, i.e., $row_{i}(\mathcal{N}^{^{\prime}})=\sum\limits_{j=1}^{s}a_{ij}row_{j}(\mathcal{N})$, where $(a_{i1},\dots,a_{is})^{\top}=a\in S_{a}$.
    2. (ii)
      Assume that at least two columns in $\mathcal{N}$ are the same. Without loss of generality, assume that the first $m_{1}$ columns of $\mathcal{N}$ are different and for any $i\in\{m_{1}+1,\dots,m\}$, there exists $j\in\{1,\dots,m_{1}\}$ such that $col_{i}(\mathcal{N})=col_{j}(\mathcal{N})$. By (21), for any $i\in\{1,\dots,s\}$, there exists $(a_{i1}^{^{\prime}},\dots,a_{is}^{^{\prime}})^{\top}\in\mathbb{Q}^{s}$ such that
      $row_{i}(\mathcal{N}^{^{\prime}})=\sum\limits_{j=1}^{s}a_{ij}^{^{\prime}}row_{j}(\mathcal{N}).$ (28)
      We denote by $\mathcal{N}_{1}$ and $\mathcal{N}^{^{\prime}}_{1}$ the first $m_{1}$ columns of $\mathcal{N}$ and $\mathcal{N}^{^{\prime}}$. Then, by (28), we have for any $i\in\{1,\dots,s\}$,
      $row_{i}(\mathcal{N}^{^{\prime}}_{1})=\sum\limits_{j=1}^{s}a_{ij}^{^{\prime}}row_{j}(\mathcal{N}_{1}).$
      Note that all the columns of $\mathcal{N}_{1}$ are different. Then, by (i), for any $i\in\{1,\dots,s\}$, there exists $(a_{i1},\dots,a_{is})^{\top}\in S_{a}$ such that
      $row_{i}(\mathcal{N}^{^{\prime}}_{1})=\sum\limits_{j=1}^{s}a_{ij}row_{j}(\mathcal{N}_{1}).$ (29)
      Note that for any $k\in\{m_{1}+1,\dots,m\}$, there exists $j\in\{1,\dots,m_{1}\}$ such that $col_{k}(\mathcal{N})=col_{j}(\mathcal{N})$. Hence, by (28), the $k$-th and the $j$-th coordinates of $row_{i}(\mathcal{N}^{^{\prime}})$ are equal. So, by (29), for any $i\in\{1,\dots,s\}$, $row_{i}(\mathcal{N}^{^{\prime}})=\sum\limits_{j=1}^{s}a_{ij}row_{j}(\mathcal{N})$, where $(a_{i1},\dots,a_{is})^{\top}\in S_{a}$.
  2. Case 2.
    Assume that some columns of $\mathcal{N}$ are not in the set $E$. By (21), for any $i\in\{1,\dots,s\}$, there exists $(a_{i1}^{^{\prime}},\dots,a_{is}^{^{\prime}})^{\top}\in\mathbb{Q}^{s}$ such that
    $row_{i}(\mathcal{N}^{^{\prime}})=\sum\limits_{j=1}^{s}a_{ij}^{^{\prime}}row_{j}(\mathcal{N}).$ (30)
    Define two new matrices $\tilde{\mathcal{N}}$ and $\tilde{\mathcal{N}^{^{\prime}}}$ as follows. For any $j\in\{1,\ldots,m\}$, if $col_{j}(\mathcal{N})\in E$, then define
    $col_{j}(\tilde{\mathcal{N}}):=col_{j}(\mathcal{N}),\;\text{and}\;col_{j}(\tilde{\mathcal{N}^{^{\prime}}}):=col_{j}(\mathcal{N}^{^{\prime}}).$
    If $col_{j}(\mathcal{N})\not\in E$, then define
    $col_{j}(\tilde{\mathcal{N}}):=-col_{j}(\mathcal{N}),\;\text{and}\;col_{j}(\tilde{\mathcal{N}^{^{\prime}}}):=-col_{j}(\mathcal{N}^{^{\prime}}).$
    Then, by the definition of $E$, all the columns of $\tilde{\mathcal{N}}$ are in the set $E$. Note that by (30), for any $i\in\{1,\dots,s\}$,
    $row_{i}(\tilde{\mathcal{N}^{^{\prime}}})=\sum\limits_{j=1}^{s}a_{ij}^{^{\prime}}row_{j}(\tilde{\mathcal{N}}).$ (31)
    Then, by (31) and by (Case 1)–(ii), for any $i\in\{1,\dots,s\}$, there exists $(a_{i1},\dots,a_{is})^{\top}\in S_{a}$ such that
    $row_{i}(\tilde{\mathcal{N}^{^{\prime}}})=\sum\limits_{j=1}^{s}a_{ij}row_{j}(\tilde{\mathcal{N}}).$
    Therefore, by the definition of $\tilde{\mathcal{N}}$ and $\tilde{\mathcal{N}^{^{\prime}}}$, we have
    $row_{i}(\mathcal{N}^{^{\prime}})=\sum\limits_{j=1}^{s}a_{ij}row_{j}(\mathcal{N}),$
    where $(a_{i1},\dots,a_{is})^{\top}\in S_{a}$.
Example 5.
We illustrate how Algorithm 3 works for $s=d=2$.
  1. Step 1
    Enumerate all possible zero-one reactions with two species (only enumerate one direction in a pair of reversible reactions).
    $0\xrightarrow{\kappa_{1}}X_{1},$ $0\xrightarrow{\kappa_{2}}X_{2},$ $0\xrightarrow{\kappa_{3}}X_{1}+X_{2},$
    $X_{1}\xrightarrow{\kappa_{4}}X_{2},$ $X_{1}\xrightarrow{\kappa_{5}}X_{1}+X_{2},$ $X_{2}\xrightarrow{\kappa_{6}}X_{1}+X_{2}.$
    Put the corresponding stoichiometric vectors (only consider different vectors) into a set:
    $E=\{(1,0)^{\top},(0,1)^{\top},(1,1)^{\top},(-1,1)^{\top}\}.$ (32)
  2. Step 2
    Let $S_{a}:=\emptyset$. Note that $d=2$. Choose $2$ column vectors from (32). There are six possibilities as follows:
    $\quad\quad\{\begin{pmatrix}1&0\\ 0&1\end{pmatrix},\begin{pmatrix}1&1\\ 0&1\end{pmatrix},\begin{pmatrix}1&-1\\ 0&1\end{pmatrix},\begin{pmatrix}0&1\\ 1&1\end{pmatrix},\begin{pmatrix}0&-1\\ 1&1\end{pmatrix},\begin{pmatrix}1&-1\\ 1&1\end{pmatrix}\}.$
    For each matrix in the above set, do the following steps (we only show the details for the first matrix).
    1. Step 2.1.
      $\mathcal{A}=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}$.
    2. Step 2.2.
      By (22)–(24), we have
      $\Lambda_{1}=\Lambda_{2}=\emptyset,\Lambda_{3}=\{1,2\}.$
      Then, by (25), we have
      $F=$$\{(-1,-1)^{\top},(-1,0)^{\top},(-1,1)^{\top},(0,-1)^{\top},$
      $(0,0)^{\top},(0,1)^{\top},(1,-1)^{\top},(1,0)^{\top},(1,1)^{\top}\}.$
      For any $b\in F$, solve $a$ from $\mathcal{A}^{\top}a=b$. Then, we get the set of the solutions is
      $\{-1,0,1\}\times\{-1,0,1\}.$
      Add all these solutions into the set $S_{a}$.
    Similarly, we do Steps 2.1 and 2.2 for all other matrices. Then, we get the set $S_{a}$:
    $\{-1,0,1\}^{2}\cup\{(2,-1)^{\top},(2,1)^{\top},(-1,2)^{\top},(1,2)^{\top},$
    $(-1,-2)^{\top},(-2,-1)^{\top},(\frac{1}{2},\frac{1}{2})^{\top},(-\frac{1}{2},\frac{1}{2})^{\top},(\frac{1}{2},-\frac{1}{2})^{\top}\}.$
5 Implementation
We have implemented the procedure in Algorithm 1. The supporting codes are available online https://github.com/YueJ13/equivalence. As a larger example, we apply the algorithm to all the three-dimensional three-species five-reaction zero-one networks.
Step 1.
We enumerate all three-species five-reaction zero-one networks. There are $3819816$ networks.
Step 2.
We call CoefficientsForVacuous($3$, $3$), and by the computation carried out in Example 4, we get the set of rational vectors $S$ in (Step 2.). For each network from the $3819816$ networks, we do the following steps.
Step 2.1.
If the stoichiometirc matrix $\mathcal{N}$ of the network is rank three (i.e., the network is three-dimensional), and all rows of $\mathcal{N}$ are non-zero and change signs, then pick the network.
Step 2.2.
If the set $\mathcal{F}^{+}(\mathcal{N})$ is not empty, then pick the network.
Step 2.3.
If $c=3$ or $4$, then delete the network. (Note here, $c$ denotes the number of zero columns of the reactant matrix $\mathcal{Y}$.)
Step 2.4.
If there exists $a\in S$ such that
  • (1)
    $\mathcal{N}^{\top}a\neq\vec{0}$, and
  • (2)
    for any $i\in\{1,2,3\}$, if the $i$-th coordinate of $\mathcal{N}^{\top}a$ is non-zero, then $col_{i}(\mathcal{Y})=\vec{0}$,
then delete the network.
Then, we get $484477$ networks.
Step 3.
Delete the same networks after relabeling the species. Then, we get $76752$ networks.
Step 4.
We call CoefficientsForEquivalence($3$, $3$), and we get a set of vectors $S_{a}\subset{\mathbb{Q}}^{3}$, where for any $a=(a_{1},a_{2},a_{3})\in S_{a}$, we have
$a_{i}\in\{\pm 4,\pm 3,\pm 2,\pm\frac{3}{2},\pm\frac{4}{3},\pm 1,\pm\frac{3}{4},\pm\frac{2}{3},\pm\frac{1}{2},\pm\frac{1}{3},\pm\frac{1}{4},0\}.$
For each pair of networks $G$ and $G^{^{\prime}}$ from the $76752$ networks in Step 3, check if $\mathcal{N}$ and $\mathcal{N}^{^{\prime}}$ (after permutation according to the description in Step 8.1 of Algorithm 1) satisfy the following condition: for any $i\in\{1,2,3\}$, there exist $(a_{1},a_{2},a_{3})^{\top}\in S_{a}$ and $(b_{1},b_{2},b_{3})^{\top}\in S_{a}$ such that
$\qquad\quad\enspace row_{i}(\mathcal{N}^{^{\prime}})=\sum\limits_{j=1}^{3}a_{j}row_{j}(\mathcal{N}),\;row_{i}(\mathcal{N})=\sum\limits_{j=1}^{3}b_{j}row_{j}(\mathcal{N}^{^{\prime}}).$
If yes, then put the networks $G$ and $G^{^{\prime}}$ into the same class. Then, we get $40522$ classes of networks. Choosing the first network from every class, we get $40522$ networks.
Step 5.
Calculate the reduced Gröbner bases of the $40522$ networks. Delete the networks whose reduced Gröbner bases are $1$ or have a monomial as an element. Then, we get $32668$ networks.
Step 6.
Combine the networks having the same reduced Gröbner basis into the same class. Choosing the first network from every class, we get $32394$ networks.
We record the timings for carrying out the above steps in the following Table 1. In the experiments, we use the command Basis [13] in Maple [22] to calculate the reduced Gröbner bases. Note that if we do not apply Algorithm 1, then we need about $60$ hours to check the reduced Gröbner bases of all three-dimensional three-species five-reaction zero-one networks. Our purpose is to show the improvement Algorithm 1 makes comparing with purely computing Gröbner bases. One can improve the computational timings presented in Table 1 by applying any more advanced tools for computing Gröbner bases rather than the standard Maple package.
Table 1: Computational time (min: minutes; s: seconds)
STEP NUMBER OF NETWORKS TIME
Step 1 3819816 1 min
Step 2 484477 30 min
Step 3 76752 145 s
Step 4 40522 40 min
Step 5 32668 100 s
Step 6 32394 30 min
Notes: (i) The column “STEP” lists the steps in the procedure shown in Algorithm 1. (ii) The middle column records the number of networks we get after carrying out each step. (iii) The column “TIME” records the computational time for each step.
We run the procedure by a 3.60 GHz Inter Core i9 processor (64GB total memory) under Windows 10.
Acknowledgments We thank Professor Balázs Boros for his suggestion on how to efficiently generate small networks in the Joint Annual Meeting KSMB-SMB 2024. We thank Professor Murad Banaji for generating and providing all three-dimensional three-species five-reaction zero-one networks by his own efficient procedure as a reference for us.

### v2

@@
-An Efficient Algorithm for Determining the Equivalence of Zero-one Reaction Networks
+Determining the Equivalence of Small Zero-one Reaction Networks
 Abstract
-Zero-one reaction networks play a crucial role in cell signaling.
-Determining the equivalence of reaction networks is a fundamental computational problem in the field of chemical reaction networks.
-In this work, we develop an efficient method for determining the equivalence of zero-one networks.
-The efficiency comes from several criteria for determining the equivalence of the steady-state ideals arising from zero-one networks, which helps for cutting down the expenses on computing Gröbner bases.
-Experiments show that our method can successfully classify over three million networks according to their equivalence in a reasonable time.
+Zero-one reaction networks are pivotal to cellular signaling, and establishing the equivalence of such networks represents a foundational computational challenge in the realm of chemical reaction network research.
+Herein, we propose a high-efficiency approach for identifying the equivalence of zero-one networks.
+Its efficiency stems from a set of criteria tailored to judge the equivalence of steady-state ideals derived from zero-one networks, which effectively reduces the computational cost associated with Gröbner basis calculations.
+Experimental results demonstrate that our proposed method can successfully categorize more than three million networks by their equivalence within a feasible timeframe.
+Also, our computational results for two important classes of quadratic zero-one networks ($3$-dimensional with $3$ species, $6$ reactions; $4$-dimensional with $4$ species, $5$ reactions) show that they have no positive steady states for a generic choice of rate constants, implying these small networks generically exhibit neither multistability nor periodic orbits.
 keywords:
@@
 1 Introduction
-For the dynamical systems that arise from biochemical reaction networks, the following computational question is fundamental.
+The dynamical systems generated by biochemical reaction networks give rise to the following fundamental computational question.
 Question 1.
 Which reaction networks have the same steady-state ideal?
-The equivalence of steady-state ideals is one of the equivalence problem in the field of chemical reaction networks.
-The dynamical features of chemical reaction networks such as multistability, Hopf-bifurcation, and absolute concentration robustness (ACR) are linked to switch-like behavior, decision-making process, oscillation, and so on in cellular signaling [7, 23, 27, 9, 3, 25].
-Usually, looking at the steady states is the first step for studying the dynamical behaviors of a chemical reaction system arising under mass-action kinetics.
-However, it is possible for two different networks to have the same steady-state ideal (notice that the steady states can be considered as real points located in the algebraic variety generated by the steady-state ideal).
-For instance, consider the two reaction networks:
-$X_{1}+X_{2}\xrightarrow{\kappa_{1}}0,$ $X_{2}\xrightarrow{\kappa_{2}}X_{1}+X_{2},$ $0\xrightarrow{\kappa_{3}}X_{1}+X_{2},$ $0\xrightarrow{\kappa_{4}}X_{2};$ (1)
-$X_{1}+X_{2}\xrightarrow{\kappa_{1}}X_{2},$ $X_{2}\xrightarrow{\kappa_{2}}X_{1},$ $0\xrightarrow{\kappa_{3}}X_{1},$ $0\xrightarrow{\kappa_{4}}X_{2}.$ (2)
-They respectively have the following steady-state systems (see how to write down the system for a given network later in (8))
-$\{-\kappa_{1}x_{1}x_{2}+\kappa_{2}x_{2}+\kappa_{3},-\kappa_{1}x_{1}x_{2}+\kappa_{3}+\kappa_{4}\};$ (3)
-$\{-\kappa_{1}x_{1}x_{2}+\kappa_{2}x_{2}+\kappa_{3},-\kappa_{2}x_{2}+\kappa_{4}\}.$ (4)
-It is straightforward to compute the reduced Gröbner bases [8] for the above two steady-state systems, and we find that in $\mathbb{Q}(\kappa)[x]$ they have the same reduced Gröbner basis
+The congruence of steady-state ideals is one of the core equivalence problems in the field of chemical reaction networks.
+Numerous key dynamical characteristics of chemical reaction networks—including multistability, Hopf bifurcations, and absolute concentration robustness (ACR)—are correlated with switch-like behaviors, decision-making processes, oscillations, and other key events in cellular signaling [8, 22, 26, 10, 3, 24].
+Typically, examining the steady states constitutes the primary step in investigating the dynamical behaviors of a chemical reaction system arising under mass-action kinetics.
+Notice that the steady states can be considered as positive real points located in the algebraic variety generated by the steady-state ideal.
+However, it is possible for two distinct networks to have exactly the same steady-state ideal.
+For instance, consider the following two reaction networks (1) and (2).
+$X_{1}+X_{2}\xrightarrow{\kappa_{1}}0,$ $X_{2}\xrightarrow{\kappa_{2}}X_{1}+X_{2},$ $0\xrightarrow{\kappa_{3}}X_{1}+X_{2},$ $0\xrightarrow{\kappa_{4}}X_{2}$ (1)
+$X_{1}+X_{2}\xrightarrow{\kappa_{1}}X_{2},$ $X_{2}\xrightarrow{\kappa_{2}}X_{1},$ $0\xrightarrow{\kappa_{3}}X_{1},$ $0\xrightarrow{\kappa_{4}}X_{2}$ (2)
+Their steady-state ideals are respectively generated by the following two sets of polynomials (3) and (4) (we will explain how to write down these polynomials for the given networks later in (8)).
+$\{-\kappa_{1}x_{1}x_{2}+\kappa_{2}x_{2}+\kappa_{3},-\kappa_{1}x_{1}x_{2}+\kappa_{3}+\kappa_{4}\}$ (3)
+$\{-\kappa_{1}x_{1}x_{2}+\kappa_{2}x_{2}+\kappa_{3},-\kappa_{2}x_{2}+\kappa_{4}\}$ (4)
+It is straightforward to compute that the above two steady-state systems have the same reduced Gröbner basis [9] in $\mathbb{Q}(\kappa)[x]$:
 $\{\kappa_{2}x_{2}-\kappa_{4},\kappa_{1}\kappa_{4}x_{1}-\kappa_{2}\kappa_{3}-\kappa_{2}\kappa_{4}\}.$ (5)
-In many cases, the above two networks are considered to be equivalent because when we only care about the dynamical behaviors determined by the steady states such as multistationarity, multistability, Hopf-bifurcation and ACR, we only need to study one of the networks that have the same steady-state ideal.
-Notice that relabeling the species or the reactions in a network naturally gives another kind of equivalence (see more details in [5] on the “dynamical equivalence” of networks).
-For instance, if we relabel the species $X_{1}$ and $X_{2}$ and we relabel the first and the second reactions in the network (1), then we get the following network
-[…]
-Of course, the network (6) should be considered to be equivalent to the network (1) and so, it is also equivalent to the network (2).
-Hence, in our context, we say two networks are equivalent if they have the same steady-state ideal after relabeling (see Definition 1).
-Determining the equivalence is computationally challenging since a standard approach is to compute the reduced Gröbner bases.
-In a recent related work [12], the authors point out that there are few works on looking at the equivalence of the steady-state ideals, and we see that deriving explicit conditions for the equivalence is difficult in general.
-So, we might need to focus on a special class of networks for making this problem more feasible.
-The standard way (computing Gröbner bases [17, 15, 13, 14, 1, 16]) might be efficient for small networks such as one-species networks or two-reaction networks.
-However, in the studying of small networks, we usually need to deal with a substantial amount of networks.
-The idea of studying the small networks is motivated by the fact that many important dynamical features have inheritance from a large network to a smaller “subnetwork” such as multistability [20, 6], oscillation [2] and local bifurcations [4].
-In many recent works, people are searching for the smallest networks from a big class of networks that admit these dynamical behaviors.
-In these studies, we need to first determine the equivalence of networks for making the searching process easier.
-For instance, Banaji and Boros [3] recently completely classify the smallest at-most-bimolecular networks that admit Hopf bifurcations (they contain three species and four reactions).
-Tang and Wang [27] find that the smallest zero-one networks that admit Hopf bifurcations are four-dimensional (they contain four species and five reactions).
-Kaihnsa, Nguyen and Shiu [21] prove that an at-most-bimolecular network admitting both multistationarity and ACR has at least three species and three reactions, and it is at least two-dimensional.
-The focus of this work is the zero-one networks since many biological significant networks in cell signaling are zero-one such as phosphorylation-dephosphorylation cycle [18], cell cycle [24], hybrid histidine kinase [26], and so on.
-In this work, our main contribution is an efficient algorithm (Algorithm 1) for determining the equivalence of zero-one networks.
+In numerous instances, the two networks above are regarded equivalent if our focus is solely on dynamical behaviors governed by steady states, including multistationarity, multistability, Hopf bifurcations, and ACR.
+Notice that relabeling the species or the reactions in a network naturally gives another kind of equivalence, say natural equivalence.
+For example, if we relabel the species $X_{1}$ and $X_{2}$ and we relabel the first and the second reactions in network (1), then we get the following network
+$X_{1}\xrightarrow{\kappa_{1}}X_{1}+X_{2},\;\;X_{1}+X_{2}\xrightarrow{\kappa_{2}}0,\;\;0\xrightarrow{\kappa_{3}}X_{1}+X_{2},\;\;0\xrightarrow{\kappa_{4}}X_{1}.$ (6)
+Of course, network (6) should be considered equivalent to network (1).
+So, it is also equivalent to network (2).
+Hence, in our context, we say that two networks are equivalent if they have the same steady-state ideal after undergoing relabeling, see Definition 3 (for further details on the “dynamical equivalence” of networks, refer to [5]).
+Determining equivalence among a large class of reaction networks is computationally challenging, as the standard approach involves computing their respective reduced Gröbner bases.
+In a recent related work [13], the authors point out that there are too few works investigating the equivalence of steady-state ideals, and we also see that deriving explicit conditions for the equivalence is difficult in general.
+In this work, we propose focusing on a specific class of small networks to make this problem more feasible.
+The idea of studying small networks is motivated by the fact that many important dynamical features are inherited from a large network to a smaller “subnetwork” such as multistability [18, 7], oscillation [2] and local bifurcations [6].
+A growing body of recent research focuses on finding the smallest networks from extensive classes of networks that admit these dynamical behaviors, and determining network equivalence first is essential to simplify the search process [3, 26, 19, 16].
+Although computing the reduced Gröbner basis [1] for a single small network might not be difficult, determining equivalence in the aforementioned studies remains challenging, as one typically needs to handle a substantial amount of networks.
+Another core proposal of this work is to focus specifically on zero-one networks for equivalence determination.
+Zero-one networks are crucial for intracellular signal transduction primarily such as the phosphorylation cycle [15], cell cycle [23], hybrid histidine kinase [25], and so on.
+They inherently ensure signal robustness, enabling stable responses against minor internal or external fluctuations instead of ambiguous intermediate outputs, which is essential for cells to make accurate functional decisions (e.g., proliferation, apoptosis).
+In this work, our main contribution is an efficient algorithm (Algorithm 1 with sub-algorithms Algorithms 23) for determining the equivalence of zero-one networks that admit positive steady states.
 The efficiency of the proposed new method comes from several criteria for determining the equivalence of the steady-state ideals arising from zero-one networks, which dramatically cuts down the expenses on computing Gröbner bases since many computations are avoided.
-We have successfully applied the new method to determining the equivalence of the smallest zero-one networks that admit multistationarity (these networks are three-dimensional, three-species and five-reaction [19]).
-The input contains over three millions such networks, and the new algorithm returns $32394$ networks after determining the equivalence.
-The computation is completed in a short time (see Table 1).
+Note that zero-one networks are enumerable when the numbers of species and reactions are given.
+We make use of this fact to generate the possible rational vectors “aa” stated in Theorem 1, enabling efficient verification of whether a network admits positive steady states, see Algorithm 2.
+By a similar idea, we generate the possible rational vectors “aa” stated in condition (I) of Theorem 2 to efficiently check if a steady-state ideal is vacuous, see Algorithm 3.
+In some special cases, the “zero-one” assumption can be used to develop simple criteria for determining the vacuous ideals, see Theorem 3.
+Consequently, our proposed method obviates the necessity of Gröbner basis computations, and accurately excludes both networks that do not admit positive steady states and those with vacuous ideals, thus enabling highly efficient, precise pre-screening of target networks.
+This paper reports improved and extended research work relative to the preliminary version published at ISSAC 25, with specific advancements as follows.
+First, in terms of algorithm design and implementation, we make the following two main improvements.
+  • 1.
+    We have refined the first three steps of the main algorithm [16, Algorithm 1].
+We enumerate stoichiometric matrices instead of reaction networks, motivated by two key observations: a single stoichiometric matrix typically maps to numerous distinct networks, and preliminary steps including consistency verification are only necessary for stoichiometric matrices.
+  • 2.
+    We develop Theorem 1 and Algorithm 2 to verify consistency, given that most input networks do not admit positive steady states and we aim to filter them out prior to determining equivalence.
+We have retested the new algorithm to determine the equivalence of the smallest zero-one networks that admit multistationarity (three-dimensional, three-species and five-reaction [17]).
+As reported in [16], this class contains over three millions networks, and the new algorithm returns $38185$ networks after determining the equivalence.
+Experimental results confirm the new algorithm achieves $50\%$ higher efficiency, halving computational time relative to the old version.
+A comparison between [16, Table 1] and Table 1 reveals a reduction in total computational time from $1$ hour to less than $0.5$ hours.
+It is seen from Table 1 that although the total number of tested networks exceeds three millions, the number of corresponding stoichiometric matrices is merely $43520$.
+After implementing the first three steps, the number is further reduced to $2548$ as the matrices that admit no positive steady states are excluded and the natural equivalence is checked.
+At this stage, the corresponding network count drops to around $100,000$.
+Notably, the first three steps take only $3$ minutes to complete, while the old version took roughly half an hour to perform these preliminary operations.
+Second, from an application standpoint, we further apply our algorithm to two important families of small zero-one networks: three-dimensional systems (three species with six reactions) and four-dimensional systems (four species with five reactions).
+Our overarching goal is to identify all minimal zero-one networks that admit multistability [17] and Hopf bifurcations [26] from these two families, respectively.
+Given the substantial size of each family, our current testing efforts have focused exclusively on quadratic networks.
+Computational results show that no networks remain in either family after excluding inconsistent networks, those with vacuous steady-state ideals, and those with monomial steady-state ideals (see Tables 23).
+This implies that for generic choices of rate constants, such small quadratic zero-one networks lack positive steady states and thus cannot display nontrivial dynamical behaviors (e.g., multistability and periodic orbits).
+Notably, this finding aligns with the established result for two-dimensional quadratic networks [6, Lemma 3], suggesting that dynamically nontrivial small zero-one networks must be at least cubic.
 The rest of this paper is organized as follows.
-In Section 2, we review the basic concepts on reaction networks and the dynamical systems arising from mass-action kinetics.
-Also, we formally define the steady-state ideal and the equivalence of networks in our context.
-In Section 3, we prove a series of sufficient conditions (Theorems 13) for determining the equivalence of networks.
-In Section 4, we present the main algorithm (Algorithm 1) with two sub-algorithms (Algorithms 23) for determining the equivalence of zero-one networks.
-In Section 5, we illustrate how the main algorithm performs by applying it to the three-dimensional zero-one networks with three species and five reactions.
-Also, we explain the implementation details and present the computational timings.
+Section 2 reviews core concepts of reaction networks and the dynamical systems derived from mass-action kinetics, while formally defining the steady-state ideal and network equivalence within our framework.
+Section 3 establishes a suite of practical criteria—including Theorems 13 and Lemma 1—for determining network equivalence.
+Section 4 presents the main algorithm (Algorithm 1) alongside its auxiliary sub-algorithms (Algorithms 23) tailored to verifying the equivalence of zero-one networks.
+Section 5 demonstrates the main algorithm’s performance via applications to three-dimensional zero-one networks (three species, five reactions), as well as two families of quadratic zero-one networks: three-dimensional (three species, six reactions) and four-dimensional (four species, five reactions).
+This section further elaborates on implementation details and reports computational runtime results.
 2 Background
@@
 We say a network (7) is a zero-one network if it only has zero-one reactions.
+We say a reaction is quadratic, if the sum of stoichiometric coefficients $\alpha_{ij}$ in (7) equals to $2$.
+We say a network (7) is a quadratic network if it only has quadratic reactions.
 For any reaction defined in (7), the stoichiometric vector is $(\beta_{1j}-\alpha_{1j},\dots,\beta_{sj}-\alpha_{sj})^{\top}\in\mathbb{Q}^{s}$.
@@
 Under the assumption of mass-action kinetics, we describe how these concentrations change in time by the following system of ODEs:
-$\dot{x}\;=\;f(\kappa,x)\;:=\;\mathcal{N}v(\kappa,x)\;=\;\mathcal{N}\begin{pmatrix}\kappa_{1}\prod\limits_{i=1}^{s}x_{i}^{\alpha_{i1}}\\ \kappa_{2}\prod\limits_{i=1}^{s}x_{i}^{\alpha_{i2}}\\ \vdots\\ \kappa_{m}\prod\limits_{i=1}^{s}x_{i}^{\alpha_{im}}\\ \end{pmatrix},$ (8)
+$\dot{x}\;=\;f(\kappa,x)\;:=\;\mathcal{N}v(\kappa,x),$ (8)
 where $x=(x_{1},x_{2},\ldots,x_{s})^{\top}$, $v(\kappa,x)=(v_{1}(\kappa,x),\dots,v_{m}(\kappa,x))^{\top}$ and
 $v_{j}(\kappa,x)\;:=\;\kappa_{j}\prod\limits_{i=1}^{s}x_{i}^{\alpha_{ij}}.$ (9)
-By viewing the rate constants as a real vector of parameters $\kappa:=(\kappa_{1},\kappa_{2},\dots,\kappa_{m})^{\top}$, we have polynomials $f_{i}(\kappa,x)\in\mathbb{Q}(\kappa)[x]$, for $i\in\{1,\dots,s\}$.
+Treating the rate constants as a real vector of parameters $\kappa:=(\kappa_{1},\dots,\kappa_{m})^{\top}$ yields polynomials $f_{i}(\kappa,x)\in\mathbb{Q}(\kappa)[x]$, for $i\in\{1,\dots,s\}$.
 The polynomial ideal generated by $f$ in $\mathbb{Q}(\kappa)[x]$, denoted by $\langle f\rangle$, is called the steady-state ideal.
 Definition 1.
-Two reaction networks are equivalent if they have the same steady-state ideal in the sense of relabeling.
+For two $s\times m$ matrices $\mathcal{A}$ and $\hat{\mathcal{A}}$, we say the matrix $\hat{\mathcal{A}}$ has the form of the matrix $\mathcal{A}$ if we can obtain $\hat{\mathcal{A}}$ from $\mathcal{A}$ by permuting the rows and columns.
+Definition 2.
+For two networks $G$ and $\hat{G}$, we say the network $\hat{G}$ has the form of $G$ if we can obtain $\hat{G}$ from $G$ by relabeling the species $X_{1},\ldots,X_{s}$ as $\hat{X_{1}},\ldots,\hat{X_{s}}$, or by relabeling the reactions ${\mathcal{R}}_{1},\ldots,{\mathcal{R}}_{m}$ as $\hat{{\mathcal{R}}}_{1},\ldots,\hat{{\mathcal{R}}}_{m}$.
+Definition 3.
+Two networks $G_{1}$ and $G_{2}$ are equivalent if there exist $\hat{G_{1}}$ and $\hat{G_{2}}$ such that $\hat{G_{i}}$ has the form of $G_{i}$ for $i\in\{1,2\}$, and $\hat{G_{1}}$ and $\hat{G_{2}}$ have the same steady-state ideal.
 For any given rate-constant vector $\kappa^{*}\in{\mathbb{R}}^{m}_{>0}$, a steady state of (8) is a vector of concentrations $x^{*}\in\mathbb{R}_{\geq 0}^{s}$ such that $f(\kappa^{*},x^{*})=\vec{0}$, where $f(\kappa,x)$ is on the right-hand side of the ODEs (8), and $\vec{0}$ denotes the column vector whose coordinates are all zero.
 If all the coordinates of a steady state $x^{*}$ are strictly positive (i.e., $x^{*}\in\mathbb{R}_{>0}^{s}$), then we call $x^{*}$ a positive steady state.
+In applications, we usually only care about positive steady states.
+We say a network admits positive steady states (or, consistent) if there exists a rate-constant vector such that system (8) has at least one positive steady state.
 For the stoichiometric matrix $\mathcal{N}\in\mathbb{Q}^{s\times m}$, the positive flux cone of $\mathcal{N}$ is defined as
 $\mathcal{F}^{+}(\mathcal{N})\;:=\;\{\gamma\in{\mathbb{R}}_{>0}^{m}|\mathcal{N}\gamma=\vec{0}\}.$ (10)
-In applications, we usually only care about positive steady states of a network.
 It is well-known that a network admits positive steady states if and only if $\mathcal{F}^{+}(\mathcal{N})\neq\emptyset$.
 3 Theorem
+3.1 Sufficient conditions for inconsistent
+Theorem 1.
+Consider a network $G$ with a stoichiometric matrix $\mathcal{N}$.
+If there exists $a\in\mathbb{Q}^{s}$ such that $a^{\top}\mathcal{N}$ is non-zero and does not change sign, then the network $G$ admits no positive steady states.
+Proof.
+Let $f$ be the steady-state system defined as in (8). Let $v=(v_{1},\dots,v_{m})^{\top}$ be defined as in (9). By (8), we have
+$a^{\top}\mathcal{N}v(\kappa,x)=\sum\limits_{i=1}^{s}a_{i}f_{i}(\kappa,x),$ (11)
+where $a_{i}$ denotes the coordinates of $a$.
+Assume that the network $G$ admits positive steady states.
+We prove the conclusion by deducing a contradiction.
+Note that $a^{\top}\mathcal{N}$ is non-zero and does not change sign.
+Hence, for any rate constant vector $\kappa\in\mathbb{Q}^{m}_{>0}$ and for any corresponding positive steady state $x\in\mathbb{Q}^{s}_{>0}$, the coefficients in $a^{\top}\mathcal{N}v(\kappa,x)$ are all positive or all negative.
+By (9), for any rate constant vector $\kappa\in\mathbb{Q}^{m}_{>0}$ and for any corresponding positive steady state $x\in\mathbb{Q}^{s}_{>0}$, we have $v(\kappa,x)\in\mathbb{Q}^{m}_{>0}$.
+Hence, $a^{\top}\mathcal{N}v(\kappa,x)$ is positive or negative.
+Then, by (11), for any rate constant vector $\kappa\in\mathbb{Q}^{m}_{>0}$ and for any positive steady state $x\in\mathbb{Q}^{s}_{>0}$, $\sum\limits_{i=1}^{s}a_{i}f_{i}(\kappa,x)$ is positive or negative.
+It is contrary to $x$ is a steady state.
+Hence, the network $G$ admits no positive steady states.
+∎
+Remark 1.
+We remark that the hypotheses of Theorem 1 can be computationally checked by algebraic tools such as FindInstance in Mathematica [21].
+Also note that if a matrix $\mathcal{N}$} has a non-zero row that does not change sign, then the matrix satisfies the hypotheses of Theorem 1.
+See the following example.
+Example 1.
+Consider the following zero-one network:
+$0\xrightarrow{\kappa_{1}}X_{3},\;\;0\xrightarrow{\kappa_{2}}X_{2},\;\;0\xrightarrow{\kappa_{3}}X_{2}+X_{3},\;\;X_{1}+X_{2}+X_{3}\xrightarrow{\kappa_{4}}0.$
+The stoichiometric matrix $\mathcal{N}$ is
+$\begin{pmatrix}0&0&0&-1\\ 0&1&1&-1\\ 1&0&1&-1\end{pmatrix}$.
+Let $a=(1,0,0)^{\top}$.
+Note that $a^{\top}\mathcal{N}=(0,0,0,-1)$.
+Hence, by Theorem 1, the network $G$ admits no positive steady states.
+On the other hand, note that $f_{1}=-\kappa_{4}x_{1}x_{2}x_{3}$, and so the steady-state system $f$ obviously has no positive solutions.
 3.2 Sufficient conditions for vacuous ideals
@@
 For any network $G$ (7), let $f$ be the steady-state system defined as in (8). If there exists $a\in\mathbb{Q}^{s}$ such that
-  • (1)
+  1. (I)
     $P:=\mathcal{N}^{\top}a\neq\vec{0}$,  and
-  • (2)
-    for any $i\in\{1,\dots,s\}$, if $P_{i}\neq 0$, then we have $col_{i}(\mathcal{Y})=\vec{0}$,
+  2. (II)
+    for any $i\in\{1,\dots,s\}$, $P_{i}\neq 0$ implies $col_{i}(\mathcal{Y})=\vec{0}$,
 then $\langle f\rangle=\langle 1\rangle$ in ${\mathbb{Q}}(\kappa)[x]$.
 Proof.
-Let $v=(v_{1},\dots,v_{m})^{\top}$ be defined as in (9). By the condition (1), we have
-$P^{\top}v=\sum_{i=1}^{s}a_{i}row_{i}(\mathcal{N})v.$ (11)
-Note that $P\neq\vec{0}$. We denote by $P_{i_{1}},\dots,P_{i_{n}}$ all the non-zero coordinates in $P$. So, by (11), we have
-$\sum_{i=1}^{s}a_{i}row_{i}(\mathcal{N})v=\sum_{j=1}^{n}P_{i_{j}}v_{i_{j}}.$ (12)
-By the condition (2), for any $j\in\{1,\dots,n\}$, we have $col_{i_{j}}(\mathcal{Y})=\vec{0}$. Hence, by (9), for any $j\in\{1,\dots,n\}$,
+Let $v=(v_{1},\dots,v_{m})^{\top}$ be defined as in (9). By condition (I), we have $P\neq\vec{0}$. And, we have
+$\sum_{j=1}^{n}P_{i_{j}}v_{i_{j}}=P^{\top}v=\sum_{i=1}^{s}a_{i}row_{i}(\mathcal{N})v,$ (12)
+where $P_{i_{1}},\dots,P_{i_{n}}$ denotes all the non-zero coordinates in $P$.
+By condition (II) , for any $j\in\{1,\dots,n\}$, we have $col_{i_{j}}(\mathcal{Y})=\vec{0}$.
+Hence, by (9), for any $j\in\{1,\dots,n\}$,
 $v_{i_{j}}=\kappa_{i_{j}}.$
@@
 We illustrate how Theorem 2 works by the following network:
-[…]
+$X_{1}\xrightarrow{\kappa_{1}}X_{2},\;\;X_{2}\xrightarrow{\kappa_{2}}X_{1},\;\;0\xrightarrow{\kappa_{3}}X_{2}.$
 The corresponding reactant matrix is
@@
 If a row (column) has both positive and negative elements, we say the row (column) changes signs.
+Notice that if there exists one non-zero row of a stoichiometric matrix $\mathcal{N}$ that does not change sign, then the corresponding network admits no positive steady states.
 Theorem 3.
@@
 We denote by $c$ the number of zero columns of the reactant matrix $\mathcal{Y}$.
-Assume that all the rows of the stoichiometric matrix $\mathcal{N}$ are non-zero and change signs.
+Assume that all the non-zero rows of the stoichiometric matrix $\mathcal{N}$ change signs.
 Then, we have the following statements.
-  1. 1.
+  1. (I)
     There is at least one non-zero column in $\mathcal{Y}$ (i.e., $c<m$).
-  2. 2.
-    If $c=m-1$ and $d\geq 2$, then $\langle f\rangle=\langle 1\rangle$ in ${\mathbb{Q}}(\kappa)[x]$.
-  3. 3.
-    If $c=m-2$ and $d=s\geq 3$, then $\langle f\rangle=\langle 1\rangle$ in ${\mathbb{Q}}(\kappa)[x]$.
-  4. 4.
+  2. (II)
+    If c=m1c=m-1 and $d\geq 2$, then $\langle f\rangle=\langle 1\rangle$ in ${\mathbb{Q}}(\kappa)[x]$.
+  3. (III)
+    If c=m2c=m-2 and $d=s\geq 3$, then $\langle f\rangle=\langle 1\rangle$ in ${\mathbb{Q}}(\kappa)[x]$.
+  4. (IV)
     If $c=0$, then $\langle f\rangle\neq\langle 1\rangle$ in ${\mathbb{Q}}(\kappa)[x]$.
 Proof.
-  1. 1.
+  1. (I)
     Recall that any stoichiometric coefficients $\alpha_{ij}$ and $\beta_{ij}$ are defined as in (7).
 By the definition of zero-one network, for any $(i,j)\in\{1,\dots,s\}\times\{1,\dots,m\}$, we have $\beta_{ij}\in\{0,1\}$.
-So, if there exists $(i,j)$ such that $\alpha_{ij}=0$, then the $(i,j)$-entry of $\mathcal{N}$ (i.e., $\beta_{ij}-\alpha_{ij}$) can only be 00 or $1$.
-Hence, if all entries of $\mathcal{Y}$ are 00 (i.e., $c=m$), then all entries of $\mathcal{N}$ are 00 or $1$.
-It is contrary to the assumption that all the rows of $\mathcal{N}$ are non-zero and change signs.
+So, if there exists $(i,j)$ such that $\alpha_{ij}=0$, then the $(i,j)$-entry of $\mathcal{N}$ (i.e., $\beta_{ij}-\alpha_{ij}$) can only be 0 or $1$.
+Hence, if all entries of $\mathcal{Y}$ are 0 (i.e., $c=m$), then every entry of $\mathcal{N}$ is 0 or $1$.
+By the definition of zero-one network, $\mathcal{N}$ is not a zero matrix.
+So, there exists a non-zero row of $\mathcal{N}$ whose entries are 0 or $1$, which is contrary to the assumption that all the non-zero rows of $\mathcal{N}$ change signs.
 Therefore, there is at least one non-zero column in $\mathcal{Y}$ (i.e., $c<m$).
-  2. 2.
-    Since all the rows of $\mathcal{N}$ are non-zero and change signs, there exists the element $-1$ in every row of $\mathcal{N}$.
+  2. (II)
+    Since all the non-zero rows of $\mathcal{N}$ change signs, there exists the element 1-1 in every non-zero row of $\mathcal{N}$.
+We define a new matrix $\mathcal{N}^{*}$ by removing the zero rows from $\mathcal{N}$.
+Also, we define $\mathcal{Y}^{*}$ by removing the same rows from $\mathcal{Y}$.
 By the definition of zero-one network, we have $\alpha_{ij}\in\{0,1\}$ and $\beta_{ij}\in\{0,1\}$.
-Hence, if there exists an $(i,j)$-entry of $\mathcal{N}$ such that the entry is $-1$ (i.e., $\beta_{ij}-\alpha_{ij}=-1$), then $\alpha_{ij}=1$.
-Hence, there exists the element $1$ in every row of $\mathcal{Y}$.
-Since the number of zero columns is $m-1$, we can assume that the first $m-1$ columns of $\mathcal{Y}$ are the zero vector.
-Note that there exists the element $1$ in every row of $\mathcal{Y}$.
-So, all the coordinates of the the $m$-th column of $\mathcal{Y}$ must be $1$.
+Hence, if there exists an $(i,j)$-entry of $\mathcal{N}^{*}$ such that the entry is 1-1 (i.e., $\beta_{ij}-\alpha_{ij}=-1$), then $\alpha_{ij}=1$.
+Hence, there exists the element $1$ in every row of $\mathcal{Y}^{*}$.
+Since the number of zero columns of $\mathcal{Y}$ is m1m-1, we can assume that the first m1m-1 columns of $\mathcal{Y}$ are the zero vectors.
+Then, by the definition of $\mathcal{Y}^{*}$, the first m1m-1 columns of $\mathcal{Y}^{*}$ are the zero vectors.
+Note that there exists the element $1$ in every row of $\mathcal{Y}^{*}$.
+So, all the coordinates of the the $m$-th column of $\mathcal{Y}^{*}$ must be $1$.
+Since $d\geq 2$ (note here, ${{\rm rank}}({\mathcal{N}})=d$), the matrix $\mathcal{Y}^{*}$ has more than one row.
 Then, we have
-    $\mathcal{Y}=\begin{pmatrix}0&\cdots&0&1\\ &\ddots&&\vdots\\ 0&\cdots&0&1\end{pmatrix}.$
-    By the definition of zero-one network, if there exists $(i,j)$ such that $\alpha_{ij}=0$, then the $(i,j)$-entry of $\mathcal{N}$ (i.e., $\beta_{ij}-\alpha_{ij}$) can only be 00 or $1$.
-Hence, every coordinate of the first $m-1$ columns of $\mathcal{N}$ can only be 00 or $1$.
-Note that there exists $-1$ in every row vector of $\mathcal{N}$.
-So, all the coordinates of the the $m$-th column of $\mathcal{N}$ are $-1$.
-Thus, $\mathcal{N}$ has the following form
-    $\mathcal{N}=\begin{pmatrix}&&&-1\\ &\mathcal{A}_{1}&&\vdots\\ &&&-1\end{pmatrix},$ (14)
-    where all the entries of the $s\times(m-1)$ submatrix $\mathcal{A}_{1}$ are 00 or $1$.
-Recall that for any $i\in\{1,\dots,s\}$, we denote by $row_{i}(\mathcal{N})$ the $i$-th row of $\mathcal{N}$.
-Since $d\geq 2$ (note here, ${\tt rank}({\mathcal{N}})=d$), there exist $i,j\in\{1,\dots,s\}$ such that $P^{\top}:=row_{i}(\mathcal{N})-row_{j}(\mathcal{N})\neq\vec{0}^{\top}$.
+    ${\mathcal{Y}^{*}}=\begin{pmatrix}0&\cdots&0&1\\ &\ddots&&\vdots\\ 0&\cdots&0&1\end{pmatrix}$.
+    By the definition of zero-one network, if there exists $(i,j)$ such that $\alpha_{ij}=0$, then the $(i,j)$-entry of $\mathcal{N}^{*}$ (i.e., $\beta_{ij}-\alpha_{ij}$) can only be 0 or $1$.
+Hence, every coordinate of the first m1m-1 columns of $\mathcal{N}^{*}$ can only be 0 or $1$.
+Note that there exists 1-1 in every row vector of $\mathcal{N}^{*}$.
+So, all the coordinates of the the $m$-th column of $\mathcal{N}^{*}$ are 1-1.
+Thus, $\mathcal{N}^{*}$ has the following form
+    ${\mathcal{N}^{*}}=\begin{pmatrix}&&&-1\\ &\mathcal{A}_{1}&&\vdots\\ &&&-1\end{pmatrix}$, (14)
+    where all the entries of the submatrix $\mathcal{A}_{1}$ are 0 or $1$.
+Recall that for any $i\in\{1,\dots,s\}$, $row_{i}(\mathcal{N})$ denotes the $i$-th row of $\mathcal{N}$.
+Since $d\geq 2$, there exist two non-zero rows $row_{i}(\mathcal{N})$ and $row_{j}(\mathcal{N})$ such that $P^{\top}:=row_{i}(\mathcal{N})-row_{j}(\mathcal{N})\neq\vec{0}^{\top}$.
 So, if $P_{k}\neq 0$, then by (14), we have $k\in\{1,\dots,m-1\}$.
@@
 Therefore, by Theorem 2, we have $\langle f\rangle=\langle 1\rangle$.
-  3. 3.
-    Assume that the first $m-2$ columns of $\mathcal{Y}$ are the zero vector. Then, by the definition of zero-one network, we have
+  3. (III)
+    Assume that the first m2m-2 columns of $\mathcal{Y}$ are the zero vector. Then, by the definition of zero-one network, we have
     $\mathcal{Y}=\begin{pmatrix}0&\cdots&0&\\ &\ddots&&\mathcal{A}_{2}\\ 0&\cdots&0&\end{pmatrix}$,
-    where the entries of the $s\times 2$ submatrix $\mathcal{A}_{2}$ are 00 or $1$. Then, similar to the proof of (2), by the definition of zero-one network, we have
+    where the entries of the $s\times 2$ submatrix $\mathcal{A}_{2}$ are 0 or $1$.
+Since $d=s$, all rows of $\mathcal{N}$ are non-zero.
+Then, similar to the proof of (2), by the definition of zero-one network, we have
     $\mathcal{N}=\begin{pmatrix}\mathcal{A}_{3}&\mathcal{A}_{4}\end{pmatrix}$, (15)
-    where the entries of the $s\times(m-2)$ submatrix $\mathcal{A}_{3}$ are 00 or $1$.
-Since all rows of $\mathcal{N}$ are non-zero and change signs, there exists the element $-1$ in every row of $\mathcal{N}$.
+    where the entries of the $s\times(m-2)$ submatrix $\mathcal{A}_{3}$ are 0 or $1$.
+Since all rows of $\mathcal{N}$ change signs, there exists the element 1-1 in every row of $\mathcal{N}$.
 Hence, the rows of the $s\times 2$ submatrix $\mathcal{A}_{4}$ must be in the set
@@
 Then, we have $P_{m-1}=P_{m}=0$.
-Since ${\tt rank}(\mathcal{N})=d=s$, we have $P\neq\vec{0}$.
+Since ${{\rm rank}}(\mathcal{N})=d=s$, we have $P\neq\vec{0}$.
 Therefore, for any $k\in\{1,\dots,m\}$, if $P_{k}\neq 0$, we have $k\in\{1,\dots,m-2\}$.
@@
 Therefore, by Theorem 2, we have $\langle f\rangle=\langle 1\rangle$.
-  4. 4.
-    Since there is no zero columns in $\mathcal{Y}$, by (9), for any $i\in\{1,\dots,m\}$, $v_{i}=\kappa_{i}\prod_{j=1}^{s}x^{\alpha_{ji}}$, where $\sum\limits_{j=1}^{s}\alpha_{ji}\geq 0$ since $\alpha_{ji}\in\{0,1\}$.
-Hence, by (8), all terms in $f$ contain the variables $x_{i}$’s.
+  4. (IV)
+    Since there is no zero columns in $\mathcal{Y}$, by (9), for any $i\in\{1,\dots,m\}$, $v_{i}=\kappa_{i}\prod_{j=1}^{s}x^{\alpha_{ji}}$, where $\sum\limits_{j=1}^{s}\alpha_{ji}>0$ since $\alpha_{ji}\in\{0,1\}$.
+Hence, by (8), for any $i\in\{1,\dots,s\}$, if $row_{i}(\mathcal{N})$ is a zero vector, then $f_{i}=0$.
+By (8), for any $i\in\{1,\dots,s\}$, if $row_{i}(\mathcal{N})$ is non-zero, then all terms in $f_{i}$ contain the variables $x_{j}$’s.
 So, in ${\mathbb{Q}}(\kappa)[x]$, $1\notin\langle f\rangle$.
@@
 We illustrate how Theorem 3 works by the following examples.
-(1) Consider the following zero-one network:
-[…]
-$0\displaystyle 0\xrightarrow{\kappa_{4}}X_{1},\;\;X_{1}+X_{2}+X_{3}\xrightarrow{\kappa_{5}}$ 0 .
+(I) Consider the following zero-one network:
+$0\xrightarrow{\kappa_{1}}X_{3},\;\;0\xrightarrow{\kappa_{2}}X_{2},\;\;0\xrightarrow{\kappa_{3}}X_{2}+X_{3},$
+$0\xrightarrow{\kappa_{4}}X_{1},\;\;X_{1}+X_{2}+X_{3}\xrightarrow{\kappa_{5}}0.$
 The stoichiometric matrix $\mathcal{N}$ is
@@
 $\begin{pmatrix}0&0&0&0&1\\ 0&0&0&0&1\\ 0&0&0&0&1\end{pmatrix}$.
-Note that there are $4$ zero columns in $\mathcal{Y}$ (i.e., $c=4=m-1$). Hence, by Theorem 2 (2), we have $\langle f\rangle=\langle 1\rangle$. Note that
+Note that there are $4$ zero columns in $\mathcal{Y}$ (i.e., c=4=m1c=4=m-1). Hence, by Theorem 3 (II), we have $\langle f\rangle=\langle 1\rangle$. Note that
 $f_{1}=$ $\kappa_{4}-\kappa_{5}x_{1}x_{2}x_{3},$
@@
 It is straightforward to check that $\frac{f_{1}-f_{2}}{\kappa_{4}-\kappa_{2}-\kappa_{3}}=1$.
-(2) Consider the following zero-one network:
-[…]
+(II) Consider the following zero-one network:
+$0\xrightarrow{\kappa_{1}}X_{3},\;\;0\xrightarrow{\kappa_{2}}X_{2},\;\;0\xrightarrow{\kappa_{3}}X_{2}+X_{3},$
 $X_{2}\xrightarrow{\kappa_{4}}X_{1},\;\;X_{1}+X_{2}+X_{3}\xrightarrow{\kappa_{5}}0.$
@@
 $\begin{pmatrix}0&0&0&0&1\\ 0&0&0&1&1\\ 0&0&0&0&1\end{pmatrix}$.
-Note that there are $3$ zero columns in $\mathcal{Y}$ (i.e., $c=3=m-2$). Hence, by Theorem 2 (3), we have $\langle f\rangle=\langle 1\rangle$. Note that
+Note that there are $3$ zero columns in $\mathcal{Y}$ (i.e., c=3=m2c=3=m-2). Hence, by Theorem 3 (III), we have $\langle f\rangle=\langle 1\rangle$. Note that
 $f_{1}=$ $\kappa_{4}x_{2}-\kappa_{5}x_{1}x_{2}x_{3},$
@@
 It is straightforward to check that $\frac{f_{1}+f_{2}-2f_{3}}{-2\kappa_{1}+\kappa_{2}-\kappa_{3}}=1$.
-(3) Consider the following zero-one network:
+(III) Consider the following zero-one network:
 $X_{1}\xrightarrow{\kappa_{1}}X_{3},\;\;X_{1}\xrightarrow{\kappa_{2}}X_{2},\;\;X_{1}\xrightarrow{\kappa_{3}}X_{2}+X_{3},$
@@
 $\begin{pmatrix}1&1&1&0&1\\ 0&0&0&1&1\\ 0&0&0&0&1\end{pmatrix}$.
-Note that there is no zero columns in $\mathcal{Y}$ (i.e., $c=0$). Hence, by Theorem 2 (4), we have $\langle f\rangle\neq\langle 1\rangle$. Note that
+Note that there is no zero columns in $\mathcal{Y}$ (i.e., $c=0$). Hence, by Theorem 3 (IV), we have $\langle f\rangle\neq\langle 1\rangle$. Note that
 $f_{1}=$ $-\kappa_{1}x_{1}-\kappa_{2}x_{1}-\kappa_{3}x_{1}+\kappa_{4}x_{2}-\kappa_{5}x_{1}x_{2}x_{3},$
@@
 It is straightforward to check that $1\notin\langle f\rangle$.
+Remark 2.
+We remark that although Theorem 2 is straightforward to prove, the hypotheses in the theorems cannot be directly checked by a computer, while Theorem 3 is explicit enough.
+In Section 4, we solve this issue by Algorithm 3.
 3.3 Sufficient conditions for equivalence
-Theorem 3.
-For any two networks $G$ and $G^{^{\prime}}$, let $\mathcal{N}$ and $\mathcal{N}^{^{\prime}}$ be the corresponding stoichiometric matrices, and let $\mathcal{Y}$ and $\mathcal{Y}^{^{\prime}}$ be the reactant matrices.
-Let $f$ and $f^{^{\prime}}$ be the steady-state systems defined as in (8).
+Lemma 1.
+For any two networks $G$ and $\hat{G}$ with $s$ species and $m$ reactions, let $\mathcal{N}$ and $\hat{\mathcal{N}}$ be the corresponding stoichiometric matrices, and let $\mathcal{Y}$ and $\hat{\mathcal{Y}}$ be the reactant matrices.
+Let $f$ and $\hat{f}$ be the steady-state systems defined as in (8).
 If
-  • (1)
-    $\mathcal{Y}=\mathcal{Y}^{^{\prime}}$, and
-  • (2)
-    for any $i\in\{1,\dots,s\}$, we have
-    $row_{i}(\mathcal{N})\in\operatorname{span}_{\mathbb{Q}}\{row_{1}(\mathcal{N}^{^{\prime}}),\dots,row_{s}(\mathcal{N}^{^{\prime}})\},$
-    and
-    $row_{i}(\mathcal{N}^{^{\prime}})\in\operatorname{span}_{\mathbb{Q}}\{row_{1}(\mathcal{N}),\dots,row_{s}(\mathcal{N})\},$
-where $\operatorname{span}_{\mathbb{Q}}\{\cdot\}$ means the rational vector space spanned by a set of rational vectors, then $\langle f\rangle=\langle f^{^{\prime}}\rangle$ in ${\mathbb{Q}}(\kappa)[x]$.
+  1. (I)
+    $\mathcal{Y}=\hat{\mathcal{Y}}$, and
+  2. (II)
+    for any $i\in\{1,\dots,s\}$, we have row_i(N) ∈span_Q {row_1(^N), …, row_s(^N)}, and row_i(^N) ∈span_Q {row_1(N), …, row_s(N)},
+where $\operatorname{span}_{\mathbb{Q}}\{\cdot\}$ means the rational vector space spanned by a set of rational vectors, then $\langle f\rangle=\langle\hat{f}\rangle$ in ${\mathbb{Q}}(\kappa)[x]$.
 Proof.
-Let $v=(v_{1},\dots,v_{m})^{\top}$ and $v^{^{\prime}}=(v_{1}^{^{\prime}},\dots,v_{m}^{^{\prime}})^{\top}$ be defined as in (9) corresponding to $G$ and $G^{\prime}$ respectively, where
-$v_{i}=\kappa_{i}\prod_{j=1}^{s}x_{j}^{\alpha_{ji}},\;\;v_{i}^{^{\prime}}=\kappa_{i}\prod_{j=1}^{s}x_{j}^{\alpha_{ji}^{^{\prime}}}.$ (17)
-By $\mathcal{Y}=\mathcal{Y}^{^{\prime}}$, for any $i\in\{1,\dots,s\}$ and for any $j\in\{1,\dots,s\}$, $\alpha_{ij}=\alpha_{ij}^{^{\prime}}$.
-Hence, by (17), we have $v=v^{^{\prime}}$.
-So, the conclusion follows from (8) and the condition (2).
+Let $v=(v_{1},\dots,v_{m})^{\top}$ and $\hat{v}=(\hat{v}_{1},\dots,\hat{v}_{m})^{\top}$ be defined as in (9) corresponding to $G$ and $\hat{G}$ respectively, where
+$v_{i}=\kappa_{i}\prod_{j=1}^{s}x_{j}^{\alpha_{ji}},\;\;\hat{v}_{i}=\kappa_{i}\prod_{j=1}^{s}x_{j}^{\hat{\alpha}_{ji}}.$ (17)
+By $\mathcal{Y}=\hat{\mathcal{Y}}$, for any $i\in\{1,\dots,s\}$ and for any $j\in\{1,\dots,s\}$, $\alpha_{ij}=\hat{\alpha}_{ij}$.
+Hence, by (17), we have $v=\hat{v}$.
+So, the conclusion follows from (8) and condition (II) .
 ∎
+Remark 3.
+If $G$ has less reactions than $\hat{G}$ does, then we can check if there exists a subnetwork of $\hat{G}$ such that the subnetwork is equivalent to $G$.
+Here, by “subnetwork”, we mean the network consists of some reactions of $\hat{G}$ and it has the same stoichiometric subspace as $\hat{G}$, where stoichiometric subspace means the column space of the stoichiometric matrix.
+By [7, 6], if a subnetwork admits multistability or Hopf bifurcations, then the original network also admits multistability or Hopf bifurcations.
 Example 4.
-Consider the following two networks $G$ and $G^{^{\prime}}$.
+Consider the following two networks $G$ and $\hat{G}$.
 $G:$ $\quad X_{1}+X_{2}\xrightarrow{\kappa_{1}}X_{2},$ $X_{2}\xrightarrow{\kappa_{2}}0,$ $0\xrightarrow{\kappa_{3}}X_{1}+X_{2},$
-$G^{^{\prime}}:$ $\quad X_{1}+X_{2}\xrightarrow{\kappa_{1}}X_{2},$ $X_{2}\xrightarrow{\kappa_{2}}X_{1},$ $0\xrightarrow{\kappa_{3}}X_{2}.$
-Note that the two networks $G$ and $G^{^{\prime}}$ have the same set of reactants.
-Hence, the reactant matrices $\mathcal{Y}$ and $\mathcal{Y}^{^{\prime}}$ are equal.
-The stoichiometric matrices $\mathcal{N}$ and $\mathcal{N}^{^{\prime}}$ are as follows.
-$\mathcal{N}=\begin{pmatrix}-1&0&1\\ 0&-1&1\end{pmatrix},\quad\mathcal{N}^{^{\prime}}=\begin{pmatrix}-1&1&0\\ 0&-1&1\end{pmatrix}.$
-Note that $row_{1}(\mathcal{N})=row_{1}(\mathcal{N}^{^{\prime}})+row_{2}(\mathcal{N}^{^{\prime}})$, $row_{2}(\mathcal{N})=row_{2}(\mathcal{N}^{^{\prime}})$ and $row_{1}(\mathcal{N}^{^{\prime}})=row_{1}(\mathcal{N})-row_{2}(\mathcal{N})$.
-So, by Theorem 3, the ideals generated by $f$ and $f^{^{\prime}}$ are equal.
+$\hat{G}$: $\quad X_{1}+X_{2}\xrightarrow{\kappa_{1}}X_{2},$ $X_{2}\xrightarrow{\kappa_{2}}X_{1},$ $0\xrightarrow{\kappa_{3}}X_{2}.$
+Note that the two networks $G$ and $\hat{G}$ have the same set of reactants.
+Hence, the reactant matrices $\mathcal{Y}$ and $\hat{\mathcal{Y}}$ are equal.
+The stoichiometric matrices $\mathcal{N}$ and $\hat{\mathcal{N}}$ are as follows.
+$\mathcal{N}=\begin{pmatrix}-1&0&1\\ 0&-1&1\end{pmatrix},\quad\hat{\mathcal{N}}=\begin{pmatrix}-1&1&0\\ 0&-1&1\end{pmatrix}$.
+Note that
+[…]
+[…]
+[…]
+So, by Lemma 1, the ideals generated by $f$ and $\hat{f}$ are equal.
 Remark 4.
-We remark that although Theorem 1 and Theorem 3 are straightforward to prove, the hypotheses in these theorems can not be directly checked by a computer, while Theorem 2 is explicit enough.
-In the next section, we will solve this issue by Algorithms 2 and 3.
+Notice that each network is uniquely determined by a pair of stoichiometric matrix $\mathcal{N}$ and reactant matrix $\mathcal{Y}$.
+However, a stoichiometric matrix $\mathcal{N}$ might correspond to many different reactant matrices.
+For example, if a network has the following stoichiometric matrix
+$\mathcal{N}=\begin{pmatrix}1&1&0&-1\\ 1&-1&-1&1\end{pmatrix}$, (18)
+then the reactant matrix can be one of the following two matrices
+${\mathcal{Y}}_{1}=\begin{pmatrix}0&0&0&1\\ 0&1&1&0\end{pmatrix},\quad{\mathcal{Y}}_{2}=\begin{pmatrix}0&0&1&1\\ 0&1&1&0\end{pmatrix}$.
+So, for two networks, if their stoichiometric matrices $\mathcal{N}$ and $\hat{\mathcal{N}}$ have different row spaces, then we do not need to compare the reactant matrices $\mathcal{Y}$ and $\hat{\mathcal{Y}}$.
+That means in practice, we can first check condition (II) of Lemma 1, and if it holds, we check condition (I).
+In Section 4, we check if condition (II) holds by checking if the reduced row echelon forms of the matrices $\mathcal{N}$ and $\hat{\mathcal{N}}$ are equal.
+See Step 4 in Algorithm 1.
 4 Algorithm
-In this section, we propose a main algorithm (Algorithm 1) to determine the equivalence of zero-one networks with two sub-algorithms (Algorithm 2 and 3).
-The goal of the main algorithm is to efficiently classify all $d$-dimensional $s$-species $m$-reaction zero-one networks according to the equivalence determined by their steady-state ideals.
-The correctness of Algorithm 1 follows from Theorem 3 and the correctness of the two sub-algorithms.
+In this section, we propose Algorithm 1, together with two sub-algorithms (Algorithm 2 and Algorithm 3), to determine the equivalence of zero-one networks.
+The goal is to efficiently classify all $d$-dimensional zero-one networks consisting of $s$ species $m$ reactions according to the equivalence determined by their steady-state ideals.
+The correctness of Algorithm 1 follows from Theorems 13, Lemma 1 and the correctness of the sub-algorithms.
 Algorithm 1.
@@
 Input.
-The number of the species $s$, the number of the reactions $m$ and the dimension $d$
+The number of species $s$, the number of reactions $m$ and the dimension $d$
 Output.
-Finitely many classes of $d$-dimensional $s$-species $m$-reaction zero-one networks such that the networks in the same class have the same steady-state ideal
+Finitely many classes of the steady-state systems $f$ corresponding to the $d$-dimensional $s$-species $m$-reaction zero-one networks such that the systems in the same class generate the same steady-state ideal
 Step 1.
-Enumerate all $d$-dimensional $s$-species $m$-reaction zero-one networks.
+Enumerate the stoichiometric coefficients $\alpha_{ij}$ and $\beta_{ij}$ for all $s$-species zero-one reactions and compute the corresponding stoichiometric vectors $\beta-\alpha$.
+Then, enumerate all possible $s\times m$ stoichiometric matrices $\mathcal{N}$ with rank $d$ whose columns are formed by these stoichiometric vectors.
 Step 2.
-Pick the networks satisfying the following two conditions simultaneously.
-  1. 1.
-    All the rows of the stoichiometric matrix $N$ are non-zero and change signs.
-  2. 2.
-    The set $\mathcal{F}^{+}(\mathcal{N})$ defined as in (10) is not empty.
-(##\##By this step, we remove the networks definitely admitting no positive steady states.)
+We call CoefficientsForInconsistent($s$, $d$, $m$), and we get a set $S_{1}\subset\mathbb{Q}^{s}$. For each stoichiometric matrix obtained in Step 1, we do the following steps.
+Step 2.1.
+If there exists $a\in S_{1}$ such that the vector $a^{\top}\mathcal{N}$ is non-zero and does not change sign, then delete this matrix.
+Step 2.2.
+If the set $\mathcal{F}^{+}(\mathcal{N})$ is empty, then delete this matrix.
+(#\#By Step 2, we remove the stoichiometric matrices whose all possible corresponding networks definitely admit no positive steady states by Theorem 1 and a standard way.)
 Step 3.
-For each remaining network, do the following steps:
-Step 3.1.
-If $c=m-1$ and $d\geq 2$, then delete the network.
-Step 3.2.
-If $c=m-2$ and $d=s\geq 3$, then delete the network.
-(##\##Note here, $c$ denotes the number of zero columns of the reactant matrix $\mathcal{Y}$.)
+For each remaining stoichiometric matrix, delete the stoichiometric matrices that have the same form with it.
 Step 4.
-We call CoefficientsForVacuous($s$, $d$), and we get a set $S\subset\mathbb{Q}^{s}$.
+For every remaining stoichiometric matrix $\mathcal{N}$, generate all possible corresponding reactant matrices $\mathcal{Y}$ for zero-one networks.
+For every pair $\{\mathcal{N},\mathcal{Y}\}$, we do the following steps.
+Step 4.1.
+If
+  1. (I)
+    c=m1c=m-1, and $d\geq 2$, or
+  2. (II)
+    c=m2c=m-2, and $d=s\geq 3$,
+then delete the pair. (#\#Note here, $c$ denotes the number of zero columns of the reactant matrix $\mathcal{Y}$.)
+Step 4.2.
+We call CoefficientsForVacuous($s$, $d$), and we get a set $S_{2}\subset\mathbb{Q}^{s}$. If there exists $a\in S_{2}$ such that
+  1. (I)
+    $\mathcal{N}^{\top}a\neq\vec{0}$, and
+  2. (II)
+    for any $i\in\{1,\dots,s\}$, the $i$-th coordinate of $\mathcal{N}^{\top}a$ is non-zero implies $col_{i}(\mathcal{Y})=\vec{0}$,
+then delete the pair.
+(#\#By Step 4, we remove the networks that have vacuous steady-state ideals according to Theorems 23.)
 Step 5.
-For each remaining network, if there exists $a\in S$ such that
-  • (1)
-    $\mathcal{N}^{\top}a\neq 0$, and
-  • (2)
-    for any $i\in\{1,\dots,s\}$, if the $i$-th coordinate of $\mathcal{N}^{\top}a$ is non-zero, then $col_{i}(\mathcal{Y})=\vec{0}$,
-then delete the network.
+Combine the remaining pairs $\{\mathcal{N},\mathcal{Y}\}$ into the same group if they have the same reduced row echelon form of $\mathcal{N}$. For each group, we do the following steps.
+Step 5.1.
+For each pair $\{\mathcal{N},\mathcal{Y}\}$ and $\{\hat{\mathcal{N}},\hat{\mathcal{Y}}\}$, we do the following steps.
+Step 5.1.1.
+Check if there exists a permutation on the rows of $\mathcal{Y}$ such that $\mathcal{Y}$ and $\hat{\mathcal{Y}}$ are equal after permuting the rows of $\mathcal{Y}$.
+Step 5.1.1.1.
+If yes, then put the pair of matrices $\{\mathcal{N},\mathcal{Y}\}$ and $\{\hat{\mathcal{N}},\hat{\mathcal{Y}}\}$ into the same group.
+Step 5.1.1.2.
+If not, then consider the next pair of matrices.
+(#\#By Step 5.1, we check the equivalence according to Lemma 1.)
+Step 5.2.
+For every group, choose the first pair as the representative element. Note that every pair of matrices $\{\mathcal{N},\mathcal{Y}\}$ uniquely determines a network.
 Step 6.
-Delete the same networks after relabeling the species.
+For every representative network obtained from Step 5, calculate the reduced Gröbner bases for steady-state ideal $\langle f\rangle$.
+Delete the systems whose corresponding reduced Gröbner bases are $\{1\}$ or have a monomial as an element.
+(#\#Note here, if the Gröbner basis contains a single monomial, then the network corresponding to the system admits no positive steady states for a generic choice of rate constants.)
 Step 7.
-We call CoefficientsForEquivalence($s$, $d$), and we get a set $S_{a}\subset\mathbb{Q}^{s}$.
-Step 8.
-For each pair of networks $G$ and $G^{^{\prime}}$ from the remaining networks in Step 6, do the following steps.
-Step 8.1.
-If there exists a permutation on the rows of $\mathcal{Y}$ and $\mathcal{Y}^{^{\prime}}$ such that $\mathcal{Y}$ and $\mathcal{Y}^{^{\prime}}$ are equal after permutating their rows in the same way, then permutate the rows of $\mathcal{N}$ and $\mathcal{N}^{^{\prime}}$ in the same way, and go to Step 8.2.
-Otherwise, consider the next pair of networks.
-Step 8.2.
-Check if $\mathcal{N}$ and $\mathcal{N}^{^{\prime}}$ satisfy the following condition: for any $i\in\{1,\dots,s\}$, there exist $(a_{i1}\dots,a_{is})^{\top}\in S_{a}$ and $(b_{i1},\dots,b_{is})^{\top}\in S_{a}$ such that
-$\qquad\quad\enspace row_{i}(\mathcal{N}^{^{\prime}})=\sum\limits_{j=1}^{s}a_{ij}row_{j}(\mathcal{N}),\;row_{i}(\mathcal{N})=\sum\limits_{j=1}^{s}b_{ij}row_{j}(\mathcal{N}^{^{\prime}}).$
-If yes, then put the networks $G$ and $G^{^{\prime}}$ into the same class. If not, then consider the next pair of networks.
-Step 9.
-For every class generated in Step 8, choose the first network as the representative element.
-Calculate the reduced Gröbner bases for these representative networks.
-Delete the networks whose reduced Gröbner base are $\{1\}$ or have a monomial as an element.
-Step 10.
 Combine the networks with the same reduced Gröbner basis into the same class. Choose the first network as the representative element. Return these representative elements.
+Remark 5.
+Note that in Step 1, we first enumerate the stoichiometric coefficients $\alpha_{ij}$ and $\beta_{ij}$, rather than the stoichiometric vectors directly.
+This approach facilitates extensions of the algorithm, particularly for the quadratic zero-one network where we require that the sum of $\alpha_{ij}$ for every reactant equals $2$.
+Remark 6.
+We remark that Step 2.1 in Algorithm 1 generalizes [16, Algorithm 1–Step 2.1], which follows from Remark 1 and the correctness of the following Algorithm 2.
+Algorithm 2.
+CoefficientsForInconsistent
+Input.
+The number of species $s$ and the number of the reactions $m$
+Output.
+A set $S\subset\mathbb{Q}^{s}$ such that for any $m$-reaction $s$-species zero-one network, if there exists $a\in\mathbb{Q}^{s}$ such that the condition of Theorem 1 holds, then there exists $\tilde{a}\in S$ satisfying the condition of Theorem 1.
+Step 1.
+Enumerate all possible $s$-species zero-one stoichiometric vectors.
+Step 2.
+Let $S=\{e_{1},\dots,e_{s}\}$, where $e_{i}\in\mathbb{Q}^{s}$ satisfies the $i$-th coordinate is $1$ and other coordinates are 0.
+Let $w:=(w_{1},\ldots,w_{s})$ be an indeterminate vector in $\mathbb{Q}_{\geq 0}^{s}$.
+Enumerate all relations among $w_{1},\ldots,w_{s}$.
+For every relation, we do the following steps.
+Step 2.1.
+For each stoichiometric vector $\gamma$ generated in Step 1, if it satisfies $w^{\top}\gamma<0$, then delete it. If not, then keep it.
+Step 2.2.
+For the remaining stoichiometric vectors in Step 2.1, enumerate all possible choices of $m$ vectors from them, forming a stoichiometric matrix $\mathcal{N}$.
+For each $\mathcal{N}$, we do the following step.
+Step 2.2.1.
+If all non-zero rows of $\mathcal{N}$ change signs, then check if there exists $q\in\mathbb{Q}^{s}$ such that $q^{\top}\mathcal{N}$ is non-zero and does not change sign.
+(#\#Here, we use FindInstance in Mathematica.) If so, add the Cartesian product $\bigotimes\limits_{i=1}^{s}\{q_{i},-q_{i}\}$ to the set $S$.
+Proof of the Correctness of Algorithm 2 For any $m$-reaction $s$-species zero-one network with the stoichiometric matrix $\mathcal{N}$, assume that there exists $a\in\mathbb{Q}^{s}$ such that the condition of Theorem 1 holds, i.e., $a^{\top}\mathcal{N}$ is non-zero and does not change sign.
+Note that if there exists $i\in\{1,\dots,s\}$ such that ${\rm row}_{i}({\mathcal{N}})$ is non-zero and does not change sign, then $e_{i}^{\top}\mathcal{N}={\rm row}_{i}({\mathcal{N}})$ is non-zero and does not change sign.
+Therefore, by Step 2, the vector $\tilde{a}=e_{i}\in S$ satisfies the condition of Theorem 1.
+Below, we assume that all non-zero rows of $\mathcal{N}$} change signs.
+  • (I)
+    Assume that all entries of $a^{\top}\mathcal{N}$ are non-negative.
+Let $\hat{a}=|a|\in\mathbb{Q}_{\geq 0}^{s}$.
+Then, there exists a relation of $w$ in Step 2 that satisfies the same relation of $\hat{a}$.
+Let $\hat{\mathcal{N}}$ be the matrix formed by the rows ${\rm sign}(a_{i}){\rm row}_{i}({\mathcal{N}})$ $(i=1,\ldots,s)$.
+Note that
+    $a^{\top}\mathcal{N}=\hat{a}^{\top}\hat{\mathcal{N}}$.
+    Since $a^{\top}\mathcal{N}$ is non-zero and all its entries are non-negative, the same holds for $\hat{a}^{\top}\hat{\mathcal{N}}$.
+Hence, for every column vector $\gamma$ of the matrix $\hat{\mathcal{N}}$, we have $\hat{a}^{\top}\gamma\geq 0$.
+Then, by Step 2.1 and Step 2.2, $\hat{\mathcal{N}}$ is one of the enumerated matrices under the relation of $\hat{a}$.
+Since all non-zero rows of $\mathcal{N}$ change signs, then every non-zero row of $\hat{\mathcal{N}}$ also changes sign.
+Then, by Step 2.2.1.2, the vector $q\in\mathbb{Q}^{S}$ obtained via FindInstance satisfies that $q^{\top}\hat{\mathcal{N}}$ is non-zero and all entries are non-negative.
+Let $\tilde{a}=({\rm sign}(a_{1})q_{1},\ldots,{\rm sign}(a_{s})q_{s})$.
+By Step 2.2.1.2, we know that $\tilde{a}\in S$.
+Notice that $\tilde{a}^{\top}\mathcal{N}=q^{\top}\hat{\mathcal{N}}$.
+Hence, $\tilde{a}^{\top}\mathcal{N}$ is non-zero and all its entries are non-negative, i.e., $\tilde{a}$ satisfies the condition of Theorem 1.
+  • (II)
+    Assume that all entries of $a^{\top}\mathcal{N}$ are non-positive.
+Notice that $a^{\top}\mathcal{N}$ is non-zero and all entries of $a^{\top}\mathcal{N}$ are non-negative.
+Then, the proof is similar to (I).
+Example 5.
+We illustrate how Algorithm 2 works for $m=5$ and $s=3$.
+Step 1.
+Enumerate all possible $3$-species stoichiometric vectors according to zero-one reactions, i.e., all vectors in
+$\{-1,0,1\}^{3}\setminus(0,0,0)^{\top}.$
+Step 2.
+Let $S=\{(0,0,1)^{\top},(0,1,0)^{\top},(0,0,1)^{\top}\}$.
+Enumerate all relations among $w_{1}$, $w_{2}$, and $w_{3}$, where $w_{i}\in\mathbb{Q}_{\geq 0}$.
+By symmetry, we only need to consider the following cases.
+$w_{1}>w_{2}>w_{3},\ w_{1}=w_{2}>w_{3},\ w_{1}>w_{2}=w_{3},\ w_{1}=w_{2}=w_{3}.$
+We treat every relation as follows.
+  • (I)
+    $w_{1}>w_{2}>w_{3}$.
+    Step 2.1.
+    For each stoichiometric vector $\gamma$ generated in Step 1, if it satisfies $w^{\top}\gamma<0$, then delete it. If not, then keep it. For example,
+    • (i)
+      we delete $(-1,-1,-1)^{\top}$ since $w^{\top}(-1,-1,-1)^{\top}=-w_{1}-w_{2}-w_{3}<0$;
+    • (ii)
+      we keep $(-1,1,1)^{\top}$ since we cannot determine the sign of $w^{\top}(-1,1,1)^{\top}=-w_{1}+w_{2}+w_{3}$ under the condition $w_{1}>w_{2}>w_{3}\geq 0$.
+    The remaining vectors are
+    $(-1,1,1)^{\top},$ $(0,0,1)^{\top},$ $(0,1,-1)^{\top},$ $(0,1,0)^{\top},$ $(0,1,1)^{\top},$
+    $(1,-1,-1)^{\top},$ $(1,-1,0)^{\top},$ $(1,-1,1)^{\top},$ $(1,0,-1)^{\top},$ $(1,0,0)^{\top},$
+    $(1,0,1)^{\top},$ $(1,1,-1)^{\top},$ $(1,1,0)^{\top},$ $(1,1,1)^{\top}.$
+    Step 2.2.
+    For the remaining stoichiometric vectors in Step 2.1, enumerate all possible choices of $5$ vectors from them, forming a stoichiometric matrix $\mathcal{N}$.
+For each $\mathcal{N}$, we do Step 2.2.1 of Algorithm 2.
+For example, consider
+    $\mathcal{N}=\begin{pmatrix}-1&0&0&0&1\\ 1&0&1&1&-1\\ 1&1&-1&0&-1\end{pmatrix}$.
+    Step 2.2.1.
+    Notice that all non-zero rows of $\mathcal{N}$ change signs.
+Using FindInstance in Mathemetica, we find $q=(2,1,1)^{\top}$ such that $q^{\top}\mathcal{N}$ is non-zero and does not change sign.
+Then, add the Cartesian product $\{2,-2\},\times\{1,-1\},\times\{1,-1\}$ to set $S$.
+    After enumerating all possible matrices, we add the follow set of vectors into $S$:
+    $\{0,\pm\frac{1}{4},\pm\frac{1}{3},\pm\frac{1}{2},\pm\frac{2}{3},\pm\frac{3}{4},\pm 1,\pm\frac{3}{2},\pm 2,\pm 3\}^{3}.$
+  • (II)
+    We implement the similar steps as case (I) for the other relations $w_{1}=w_{2}>w_{3}$, $w_{1}>w_{2}=w_{3}$ and $w_{1}=w_{2}=w_{3}$. And finally, we get
+    $S=\{0,\pm\frac{1}{4},\pm\frac{1}{3},\pm\frac{1}{2},\pm\frac{2}{3},\pm\frac{3}{4},\pm 1,\pm\frac{3}{2},\pm 2,\pm 3\}^{3}.$ (19)
 Algorithm 3.
@@
 Output.
-A set $S\subset\mathbb{Q}^{s}$ such that for any $d$-dimensional $s$-species zero-one network, if there exists $a\in\mathbb{Q}^{s}$ such that the conditions (1) and (2) of Theorem 1 hold, then there exists $\tilde{a}\in S$ satisfying the conditions (1) and (2) of Theorem 1
+A set $S\subset\mathbb{Q}^{s}$ such that for any $d$-dimensional $s$-species zero-one network, if there exists $a\in\mathbb{Q}^{s}$ such that conditions (I) and (II) of Theorem 2 hold, then there exists $\tilde{a}\in S$ satisfying conditions (I) and (II) of Theorem 2
 Step 1.
@@
 Step 2.
-Let $S:=\emptyset$. For every $n\in\{1,\dots,d-1\}$, choose $n$ linearly independent reactions from the reactions enumerated in Step 1. For every choice, do the following steps.
+Let $S:=\emptyset$.
+For every $n\in\{1,\dots,d-1\}$, for all possible choices of $n$ linearly independent reactions from the reactions enumerated in Step 1, we do the following steps.
+Note that here we operate on all possible combinations of $n$ reactions.
 Step 2.1.
@@
 Step 2.2.
-Calculate the basis of the null space of $\mathcal{M}$ and add all the vectors in the basis into the set $S$.
-(##\##Note here, all these vectors in the basis are in $\mathbb{Q}^{s}$.)
+Calculate a basis of the nullspace of $\mathcal{M}$ in $\mathbb{Q}^{s}$ and add all the vectors in the basis into the set $S$.
 Step 3.
 Return the set $S$.
-Proof of the Correctness of Algorithm 2. Define
-$\Delta_{1}$ :=$:=\{i\in\{1,\dots,m\}|col_{i}(\mathcal{Y})\neq\vec{0}\},$ (18)
-$\Delta_{2}$ :=$:=\{1,\dots,m\}\setminus\Delta_{1}.$ (19)
+Proof of the Correctness of Algorithm 3. For any $d$-dimensional $s$-species zero-one network, suppose $\mathcal{Y}$} and $\mathcal{N}$} are the reactant matrix and the stoichiometric matrix. Define
+$\Delta_{1}$ $:=\{i\in\{1,\dots,m\}|col_{i}(\mathcal{Y})\neq\vec{0}\},$ (20)
+$\Delta_{2}$ $:=\{1,\dots,m\}\setminus\Delta_{1}.$ (21)
 Let $M$ and $M_{k}$ ($k\in\{1,2\}$) be the real vector spaces spanned by the columns of $\mathcal{N}$} and the set of column vectors $\{col_{i}({\mathcal{N}})|i\in\Delta_{k}\}$, respectively.
-By (18) and the condition (2) of Theorem 1, we have
+If there exists $a\in\mathbb{Q}^{s}$ such that conditions (I) and (II) of Theorem 2 hold, then by (20) and condition (II) of Theorem 2, we have
 $a\in M_{1}^{\perp}.$
-Since $\mathcal{N}^{\top}a\neq\vec{0}$ (the condition (1) of Theorem 1), we have $a\not\in M^{\perp}$.
-Since $M_{1}\subset M$, we have ${\tt dim}(M_{1})<{\tt dim}(M)=d$.
-So, by Step 2 of Algorithm 2, there exists $B\subset S$ such that $B$ is the basis of $\mathcal{M}^{\perp}_{1}$.
+Since $\mathcal{N}^{\top}a\neq\vec{0}$ (condition (I) of Theorem 2), we have $a\not\in M^{\perp}$.
+Since $M_{1}\subset M$, we have ${{\rm dim}}(M_{1})<{{\rm dim}}(M)=d$.
+So, by Step 2 of Algorithm 3, there exists $B\subset S$ such that $B$ is a basis of $M^{\perp}_{1}$.
 Notice that there exists $\tilde{a}\in B$ such that $\tilde{a}\not\in M^{\perp}$, i.e., $\mathcal{N}^{\top}\tilde{a}\neq\vec{0}$.
 Otherwise, we have $B\subset M^{\perp}$ and hence, we have $M_{1}^{\perp}\subset M^{\perp}$, which is contrary to the fact that $a\in M_{1}^{\perp}\backslash M^{\perp}$.
-So, the conditions (1) and (2) of Theorem 1 hold for $\tilde{a}\in{\mathbb{Q}}^{s}$.
+So, conditions (I) and (II) of Theorem 2 hold for $\tilde{a}\in{\mathbb{Q}}^{s}$.
 Example 6.
 We illustrate how Algorithm 3 works for $s=d=3$.
-  1. Step 1.
+Step 1.
 Enumerate all three-species zero-one reactions (there are $56$ reactions in total).
-  2. Step 2.
+Step 2.
 Let $S:=\emptyset$.
@@
 For every $n\in\{1,2\}$, choose $n$ linearly independent reactions from the $56$ reactions enumerated in Step 1 and for every choice, do the following steps (below we only give details on the first round).
-    1. Step 2.1
+Step 2.1
 Suppose $(0,0,1)$ is the stoichiometric vector corresponding to the first reaction from the $56$ reactions. Let $\mathcal{M}:=(0,0,1)$.
-    2. Step 2.2
+Step 2.2
 The basis of the null space of $\mathcal{M}$ is $\{(1,0,0)^{\top},(0,1,0)^{\top}\}$. Add the vectors in the basis into the set $S$.
 Similarly, we do Steps 2.1 and 2.2 for all the other choices of reactions. Then, we get the set $S$:
-    $\{0,1\}^{3}\cup(\{-1,1\}^{s}\times\{2\})\cup$
-    ($(\{-1,1\}\times\{-2,2\}\times\{1\})\cup(\{-2,2\}\times\{-1,1\}\times\{1\})\cup$
+$\{0,1\}^{3}\cup(\{-1,1\}^{{2}}\times\{2\})\cup$
+$(\{-1,1\}\times\{-2,2\}\times\{1\})\cup(\{-2,2\}\times\{-1,1\}\times\{1\})\cup$
 $\{(-1,-1,1)^{\top},(1,-1,1)^{\top},(-1,1,1)^{\top},(-1,0,1)^{\top},$
 $(-1,1,0)^{\top},(0,-1,1)^{\top}\}\setminus(0,0,0)^{\top}.$ (22)
-Algorithm 3.
-CoefficientsForEquivalence
-Input.
-The number of species $s$ and the dimension $d$
-Output.
-A set $S_{a}\subset\mathbb{Q}^{s}$ such that if two $d$-dimensional $s$-species zero-one networks satisfy
-  • (1)
-    ${\mathcal{Y}}={\mathcal{Y}}^{^{\prime}}$, and
-  • (2)
-    for any $i\in\{1,\dots,s\}$,
-    $row_{i}(\mathcal{N}^{^{\prime}})$ $\in\operatorname{span}_{\mathbb{Q}}\{row_{1}(\mathcal{N}),\dots,row_{s}(\mathcal{N})\},$ (21)
-then for any $i\in\{1,\dots,s\}$, there exists $(a_{i1},\dots,a_{is})^{\top}\in S_{a}$ such that
-$row_{i}(\mathcal{N}^{^{\prime}})=\sum\limits_{j=1}^{s}a_{ij}row_{j}(\mathcal{N})$
+5 Implementation
+We have implemented the procedure outlined in Algorithm 1, with supporting code publicly available at https://github.com/YueJ13/equivalence.
+We further apply this algorithm to all three-dimensional zero-one networks consisting of three species and five reactions.
 Step 1.
-Enumerate all possible zero-one reactions with $s$ species (here, we only enumerate one direction in a pair of reversible reactions).
-Write the corresponding stoichiometric vectors into a set $E\subset\mathbb{Q}^{s}$ (here, it is possible for different reactions to have the same stoichiometric vector, and we only put different vectors into the set $E$).
+Enumerate the stoichiometric coefficients $\alpha_{ij}$ and $\beta_{ij}$ for all $3$-species zero-one reactions and compute the corresponding stoichiometric vectors $\beta-\alpha$.
+Then enumerate all possible $3\times 5$ stoichiometric matrices $\mathcal{N}$ of rank $3$ whose columns are stoichiometric vectors.
+There are $43520$ matrices.
 Step 2.
-Let $S_{a}:=\emptyset$. For any $d$ column vectors from $E$, do the following steps.
+We call CoefficientsForInconsistent($3$, 33, 55), and by the computation carried out in Example 5, we get a set of rational vectors $S_{1}$ in (19).
+For each stoichiometric matrix obtained in Step 1, we do the following steps.
 Step 2.1.
-Let $\mathcal{A}$ be the $s\times d$ matrix consisting of the $d$ column vectors. If $\mathcal{A}$ has any zero rows, then stop and go to the next group of $d$ vectors chosen from $E$.
+If there exists $a\in S_{1}$ such that the vector $a^{\top}\mathcal{N}$ is non-zero and does not change sign, then delete this matrix.
 Step 2.2.
-Denote by $\vec{1}$ and $-\vec{1}$ the column vectors whose coordinates are all $1$ and $-1$, respectively. Define
-$\Lambda_{1}$ :=$:=\{i\in\{1,\dots,d\}|col_{i}(\mathcal{A})=\vec{1}\},$ (22)
-$\Lambda_{2}$ :=$:=\{i\in\{1,\dots,d\}|col_{i}(\mathcal{A})=-\vec{1}\},$ (23)
-$\Lambda_{3}$ :=$:=\{1,\dots,d\}\setminus(\Lambda_{1}\cup\Lambda_{2}).$ (24)
-Define a set of vectors
-$F:=\{0,1\}^{|\Lambda_{1}|}\times\{-1,0\}^{|\Lambda_{2}|}\times\{-1,0,1\}^{|\Lambda_{3}|}\subset\mathbb{Q}^{d}.$ (25)
-(##\##Notice here $|\Lambda_{1}|+|\Lambda_{2}|+|\Lambda_{3}|=d$.)
-For any column vector $b$ in the set $F$, do the following steps.
-Step 2.2.1.
-Solve $a\in\mathbb{Q}^{s}$ from $\mathcal{A}^{\top}a=b$.
-Step 2.2.2.
-If there exists finitely many solutions, then add the solutions into the set $S_{a}$.
-If there exists infinitely many solutions, then add one of the solutions into the set $S_{a}$.
-If there is no solutions, then go to the next vector $b\in F$.
+If the set $\mathcal{F}^{+}(\mathcal{N})$ is empty, then delete this matrix.
+Then, we get $14444$ matrices.
 Step 3.
-Return the set $S_{a}$.
-Proof of the Correctness of Algorithm 3.
-Consider two $d$-dimensional $s$-species zero-one networks $G$ and $G^{^{\prime}}$.
-Assume that their reactant matrices are equal, i.e., $\mathcal{Y}=\mathcal{Y}^{^{\prime}}$.
-First, we prove a claim.
-Claim.
-If there exists $\mathcal{A}$ in Step 2.1 of Algorithm 3 such that $\mathcal{A}=\mathcal{N}$, then for any $j\in\{1,\dots,s\}$, there exists $b\in F$ (defined as in (25)) such that $row_{j}(\mathcal{N}^{^{\prime}})^{\top}=b$.
-In fact, by (22), for any $i\in\Lambda_{1}$, all the coordinates of $col_{i}(\mathcal{A})$ are $1$.
-Then, since $\mathcal{N}=\mathcal{A}$, for any $i\in\Lambda_{1}$, all the coordinates of $col_{i}(\mathcal{N})$ are $1$.
-By the definition of zero-one network, all the coordinates of $col_{i}(\mathcal{Y})$ are 00.
-Then, since $\mathcal{Y}=\mathcal{Y}^{^{\prime}}$, all the coordinates of $col_{i}(\mathcal{Y}^{^{\prime}})$ are 00.
-Thus, by the definition of zero-one network again, we have all the coordinates of $col_{i}(\mathcal{N}^{^{\prime}})$ are 00 or $1$.
-Hence, for any $j\in\{1,\dots,s\}$ and for any $i\in\Lambda_{1}$, the $i$-th coordinate of $row_{j}(\mathcal{N}^{^{\prime}})$ is 00 or $1$.
-Similarly, we can derive that for any $j\in\{1,\dots,s\}$ and for any $i\in\Lambda_{2}$, the $i$-th coordinate of $row_{j}(\mathcal{N}^{^{\prime}})$ is $-1$ or 00.
-Note that by the definition of zero-one network, for any $j\in\{1,\dots,s\}$ and for any $i\in\Lambda_{3}$, the $i$-th coordinate of $row_{j}(\mathcal{N}^{^{\prime}})$ is $-1$, 00 or $1$.
-Hence, by (25), for any $j\in\{1,\dots,s\}$, $row_{j}(\mathcal{N}^{^{\prime}})^{\top}\in F$.
-So, for any $j\in\{1,\dots,s\}$, there exists $b\in F$ such that $row_{j}(\mathcal{N}^{^{\prime}})^{\top}=b$.
-Below, we prove the correctness of Algorithm 3 by discussing two cases.
-  1. Case 1.
-    Assume that all the columns of $\mathcal{N}$ are in the set $E$ generated in Step 1 of Algorithm 3.
+For each remaining stoichiometric matrix, delete the stoichiometric matrices that have the same form with it. Then, we get $2548$ matrices.
+Step 4.
+For every remaining stoichiometric matrix $\mathcal{N}$, generate all possible corresponding reactant matrices $\mathcal{Y}$. For every pair $\{\mathcal{N},\mathcal{Y}\}$, we do the following steps.
+Step 4.1.
+If $c=3$ or $4$, then delete the pair. (#\#Note here, $c$ denotes the number of zero columns of the reactant matrix $\mathcal{Y}$.)
+Step 4.2.
+We call CoefficientsForVacuous($3$, 33), and by the compotation carried out in Example 6, we get a set of rational vectors $S_{2}$ in (Step 2.). If there exists $a\in S_{2}$ such that
   1. (I)
-      Assume that all the columns of $\mathcal{N}$ are different.
-Since all the columns in $E$ are different.
-by the choice of $\mathcal{A}$ in Step 2 of Algorithm 3, there exists $\mathcal{A}$ such that $\mathcal{A}=\mathcal{N}$.
-By (21), for any $i\in\{1,\dots,s\}$, there exists $a^{^{\prime}}:=(a_{i1}^{^{\prime}},\dots,a_{is}^{^{\prime}})^{\top}\in\mathbb{Q}^{s}$ such that
-      $row_{i}(\mathcal{N}^{^{\prime}})=\sum\limits_{j=1}^{s}a_{ij}^{^{\prime}}row_{j}(\mathcal{N}).$ (26)
-      Hence, by (26), for any $i\in\{1,\dots,s\}$, we have
-      $\mathcal{N}^{\top}a^{^{\prime}}=row_{i}(\mathcal{N}^{^{\prime}})^{\top}.$ (27)
-      By $\mathcal{A}=\mathcal{N}$ and by the claim, for any $i\in\{1,\dots,s\}$, there exists $b\in F$ such that $row_{i}(\mathcal{N}^{^{\prime}})^{\top}=b$.
-Hence, by $\mathcal{N}=\mathcal{A}$ and (27), we have $\mathcal{A}^{\top}a^{^{\prime}}=b$, i.e., there exists a solution to the equation $\mathcal{A}^{\top}a=b$.
-By Step 2.2.2, there exists $a\in S_{a}$ such that $\mathcal{A}^{\top}a=b$.
-Hence, for any $i\in\{1,\dots,s\}$, by $\mathcal{A}=\mathcal{N}$ and $row_{i}(\mathcal{N}^{^{\prime}})^{\top}=b$, we have $\mathcal{N}^{\top}a=row_{i}(\mathcal{N}^{^{\prime}})^{\top}$, i.e., $row_{i}(\mathcal{N}^{^{\prime}})=\sum\limits_{j=1}^{s}a_{ij}row_{j}(\mathcal{N})$, where $(a_{i1},\dots,a_{is})^{\top}=a\in S_{a}$.
+    $\mathcal{N}^{\top}a\neq\vec{0}$, and
   2. (II)
-      Assume that at least two columns in $\mathcal{N}$ are the same.
-Without loss of generality, assume that the first $m_{1}$ columns of $\mathcal{N}$ are different and for any $i\in\{m_{1}+1,\dots,m\}$, there exists $j\in\{1,\dots,m_{1}\}$ such that $col_{i}(\mathcal{N})=col_{j}(\mathcal{N})$.
-By (21), for any $i\in\{1,\dots,s\}$, there exists $(a_{i1}^{^{\prime}},\dots,a_{is}^{^{\prime}})^{\top}\in\mathbb{Q}^{s}$ such that
-      $row_{i}(\mathcal{N}^{^{\prime}})=\sum\limits_{j=1}^{s}a_{ij}^{^{\prime}}row_{j}(\mathcal{N}).$ (28)
-      We denote by $\mathcal{N}_{1}$ and $\mathcal{N}^{^{\prime}}_{1}$ the first $m_{1}$ columns of $\mathcal{N}$ and $\mathcal{N}^{^{\prime}}$. Then, by (28), we have for any $i\in\{1,\dots,s\}$,
-      $row_{i}(\mathcal{N}^{^{\prime}}_{1})=\sum\limits_{j=1}^{s}a_{ij}^{^{\prime}}row_{j}(\mathcal{N}_{1}).$
-      Note that all the columns of $\mathcal{N}_{1}$ are different. Then, by (i), for any $i\in\{1,\dots,s\}$, there exists $(a_{i1},\dots,a_{is})^{\top}\in S_{a}$ such that
-      $row_{i}(\mathcal{N}^{^{\prime}}_{1})=\sum\limits_{j=1}^{s}a_{ij}row_{j}(\mathcal{N}_{1}).$ (29)
-      Note that for any $k\in\{m_{1}+1,\dots,m\}$, there exists $j\in\{1,\dots,m_{1}\}$ such that $col_{k}(\mathcal{N})=col_{j}(\mathcal{N})$.
-Hence, by (28), the $k$-th and the $j$-th coordinates of $row_{i}(\mathcal{N}^{^{\prime}})$ are equal.
-So, by (29), for any $i\in\{1,\dots,s\}$, $row_{i}(\mathcal{N}^{^{\prime}})=\sum\limits_{j=1}^{s}a_{ij}row_{j}(\mathcal{N})$, where $(a_{i1},\dots,a_{is})^{\top}\in S_{a}$.
-  2. Case 2.
-    Assume that some columns of $\mathcal{N}$ are not in the set $E$.
-By (21), for any $i\in\{1,\dots,s\}$, there exists $(a_{i1}^{^{\prime}},\dots,a_{is}^{^{\prime}})^{\top}\in\mathbb{Q}^{s}$ such that
-    $row_{i}(\mathcal{N}^{^{\prime}})=\sum\limits_{j=1}^{s}a_{ij}^{^{\prime}}row_{j}(\mathcal{N}).$ (30)
-    Define two new matrices $\tilde{\mathcal{N}}$ and $\tilde{\mathcal{N}^{^{\prime}}}$ as follows. For any $j\in\{1,\ldots,m\}$, if $col_{j}(\mathcal{N})\in E$, then define
-    $col_{j}(\tilde{\mathcal{N}}):=col_{j}(\mathcal{N}),\;\text{and}\;col_{j}(\tilde{\mathcal{N}^{^{\prime}}}):=col_{j}(\mathcal{N}^{^{\prime}}).$
-    If $col_{j}(\mathcal{N})\not\in E$, then define
-    $col_{j}(\tilde{\mathcal{N}}):=-col_{j}(\mathcal{N}),\;\text{and}\;col_{j}(\tilde{\mathcal{N}^{^{\prime}}}):=-col_{j}(\mathcal{N}^{^{\prime}}).$
-    Then, by the definition of $E$, all the columns of $\tilde{\mathcal{N}}$ are in the set $E$. Note that by (30), for any $i\in\{1,\dots,s\}$,
-    $row_{i}(\tilde{\mathcal{N}^{^{\prime}}})=\sum\limits_{j=1}^{s}a_{ij}^{^{\prime}}row_{j}(\tilde{\mathcal{N}}).$ (31)
-    Then, by (31) and by (Case 1)–(ii), for any $i\in\{1,\dots,s\}$, there exists $(a_{i1},\dots,a_{is})^{\top}\in S_{a}$ such that
-    $row_{i}(\tilde{\mathcal{N}^{^{\prime}}})=\sum\limits_{j=1}^{s}a_{ij}row_{j}(\tilde{\mathcal{N}}).$
-    Therefore, by the definition of $\tilde{\mathcal{N}}$ and $\tilde{\mathcal{N}^{^{\prime}}}$, we have
-    $row_{i}(\mathcal{N}^{^{\prime}})=\sum\limits_{j=1}^{s}a_{ij}row_{j}(\mathcal{N}),$
-    where $(a_{i1},\dots,a_{is})^{\top}\in S_{a}$.
-Example 5.
-We illustrate how Algorithm 3 works for $s=d=2$.
-  1. Step 1
-    Enumerate all possible zero-one reactions with two species (only enumerate one direction in a pair of reversible reactions).
-    $0\xrightarrow{\kappa_{1}}X_{1},$ $0\xrightarrow{\kappa_{2}}X_{2},$ $0\xrightarrow{\kappa_{3}}X_{1}+X_{2},$
-    $X_{1}\xrightarrow{\kappa_{4}}X_{2},$ $X_{1}\xrightarrow{\kappa_{5}}X_{1}+X_{2},$ $X_{2}\xrightarrow{\kappa_{6}}X_{1}+X_{2}.$
-    Put the corresponding stoichiometric vectors (only consider different vectors) into a set:
-    $E=\{(1,0)^{\top},(0,1)^{\top},(1,1)^{\top},(-1,1)^{\top}\}.$ (32)
-  2. Step 2
-    Let $S_{a}:=\emptyset$. Note that $d=2$. Choose $2$ column vectors from (32). There are six possibilities as follows:
-    $\quad\quad\{\begin{pmatrix}1&0\\ 0&1\end{pmatrix},\begin{pmatrix}1&1\\ 0&1\end{pmatrix},\begin{pmatrix}1&-1\\ 0&1\end{pmatrix},\begin{pmatrix}0&1\\ 1&1\end{pmatrix},\begin{pmatrix}0&-1\\ 1&1\end{pmatrix},\begin{pmatrix}1&-1\\ 1&1\end{pmatrix}\}.$
-    For each matrix in the above set, do the following steps (we only show the details for the first matrix).
-    1. Step 2.1.
-      $\mathcal{A}=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}$.
-    2. Step 2.2.
-      By (22)–(24), we have
-      $\Lambda_{1}=\Lambda_{2}=\emptyset,\Lambda_{3}=\{1,2\}.$
-      Then, by (25), we have
-      $F=$$\{(-1,-1)^{\top},(-1,0)^{\top},(-1,1)^{\top},(0,-1)^{\top},$
-      $(0,0)^{\top},(0,1)^{\top},(1,-1)^{\top},(1,0)^{\top},(1,1)^{\top}\}.$
-      For any $b\in F$, solve $a$ from $\mathcal{A}^{\top}a=b$. Then, we get the set of the solutions is
-      $\{-1,0,1\}\times\{-1,0,1\}.$
-      Add all these solutions into the set $S_{a}$.
-    Similarly, we do Steps 2.1 and 2.2 for all other matrices. Then, we get the set $S_{a}$:
-    $\{-1,0,1\}^{2}\cup\{(2,-1)^{\top},(2,1)^{\top},(-1,2)^{\top},(1,2)^{\top},$
-    $(-1,-2)^{\top},(-2,-1)^{\top},(\frac{1}{2},\frac{1}{2})^{\top},(-\frac{1}{2},\frac{1}{2})^{\top},(\frac{1}{2},-\frac{1}{2})^{\top}\}.$
-5 Implementation
-We have implemented the procedure in Algorithm 1.
-The supporting codes are available online https://github.com/YueJ13/equivalence.
-As a larger example, we apply the algorithm to all the three-dimensional three-species five-reaction zero-one networks.
-Step 1.
-We enumerate all three-species five-reaction zero-one networks. There are $3819816$ networks.
-Step 2.
-We call CoefficientsForVacuous($3$, $3$), and by the computation carried out in Example 4, we get the set of rational vectors $S$ in (Step 2.).
-For each network from the $3819816$ networks, we do the following steps.
-Step 2.1.
-If the stoichiometirc matrix $\mathcal{N}$ of the network is rank three (i.e., the network is three-dimensional), and all rows of $\mathcal{N}$ are non-zero and change signs, then pick the network.
-Step 2.2.
-If the set $\mathcal{F}^{+}(\mathcal{N})$ is not empty, then pick the network.
-Step 2.3.
-If $c=3$ or $4$, then delete the network. (Note here, $c$ denotes the number of zero columns of the reactant matrix $\mathcal{Y}$.)
-Step 2.4.
-If there exists $a\in S$ such that
-  • (1)
-    $\mathcal{N}^{\top}a\neq\vec{0}$, and
-  • (2)
-    for any $i\in\{1,2,3\}$, if the $i$-th coordinate of $\mathcal{N}^{\top}a$ is non-zero, then $col_{i}(\mathcal{Y})=\vec{0}$,
-then delete the network.
-Then, we get $484477$ networks.
-Step 3.
-Delete the same networks after relabeling the species. Then, we get $76752$ networks.
-Step 4.
-We call CoefficientsForEquivalence($3$, $3$), and we get a set of vectors $S_{a}\subset{\mathbb{Q}}^{3}$, where for any $a=(a_{1},a_{2},a_{3})\in S_{a}$, we have
-$a_{i}\in\{\pm 4,\pm 3,\pm 2,\pm\frac{3}{2},\pm\frac{4}{3},\pm 1,\pm\frac{3}{4},\pm\frac{2}{3},\pm\frac{1}{2},\pm\frac{1}{3},\pm\frac{1}{4},0\}.$
-For each pair of networks $G$ and $G^{^{\prime}}$ from the $76752$ networks in Step 3, check if $\mathcal{N}$ and $\mathcal{N}^{^{\prime}}$ (after permutation according to the description in Step 8.1 of Algorithm 1) satisfy the following condition: for any $i\in\{1,2,3\}$, there exist $(a_{1},a_{2},a_{3})^{\top}\in S_{a}$ and $(b_{1},b_{2},b_{3})^{\top}\in S_{a}$ such that
-$\qquad\quad\enspace row_{i}(\mathcal{N}^{^{\prime}})=\sum\limits_{j=1}^{3}a_{j}row_{j}(\mathcal{N}),\;row_{i}(\mathcal{N})=\sum\limits_{j=1}^{3}b_{j}row_{j}(\mathcal{N}^{^{\prime}}).$
-If yes, then put the networks $G$ and $G^{^{\prime}}$ into the same class. Then, we get $40522$ classes of networks. Choosing the first network from every class, we get $40522$ networks.
+    for any $i\in\{1,2,3\}$, the $i$-th coordinate of $\mathcal{N}^{\top}a$ is non-zero implies $col_{i}(\mathcal{Y})=\vec{0}$,
+then delete the pair.
+Then, we get $91802$ networks.
 Step 5.
-Calculate the reduced Gröbner bases of the $40522$ networks. Delete the networks whose reduced Gröbner bases are $1$ or have a monomial as an element. Then, we get $32668$ networks.
+Combine the remaining pairs $\{\mathcal{N},\mathcal{Y}\}$ into the same group if they have the same reduced row echelon form of $\mathcal{N}$. For each group, we do the following steps.
+Step 5.1.
+For each pair $\{\mathcal{N},\mathcal{Y}\}$ and $\{\hat{\mathcal{N}},\hat{\mathcal{Y}}\}$, we do the following steps.
+Step 5.1.1.
+Check if there exists a permutation on the rows of $\mathcal{Y}$ such that $\mathcal{Y}$ and $\hat{\mathcal{Y}}$ are equal after permuting the rows of $\mathcal{Y}$.
+Step 5.1.1.1.
+If yes, then put the pair of matrices $\{\mathcal{N},\mathcal{Y}\}$ and $\{\hat{\mathcal{N}},\hat{\mathcal{Y}}\}$ into the same group.
+Step 5.1.1.2.
+If not, then consider the next pair of matrices.
+Step 5.2.
+For every group, choose the first pair $\{\mathcal{N},\mathcal{Y}\}$ as the representative element.
+Then, we get $48703$ networks (pairs).
 Step 6.
-Combine the networks having the same reduced Gröbner basis into the same class. Choosing the first network from every class, we get $32394$ networks.
-We record the timings for carrying out the above steps in the following Table 1.
+For every remaining network, calculate the reduced Gröbner bases for steady-state ideal $\langle f\rangle$.
+Delete the networks whose corresponding reduced Gröbner bases are $\{1\}$ or have a monomial as an element.
+Then, we get $38249$ networks.
+Step 7.
+Combine the networks with the same reduced Gröbner basis into the same class. Choose the first network as the representative element. Finally, we get $38185$ networks.
+We record the computational results for the above steps in Table 1.
 In the experiments, we use the command Basis [14] in Maple [20] to calculate the reduced Gröbner bases.
-Note that if we do not apply Algorithm 1, then we need about $60$ hours to check the reduced Gröbner bases of all three-dimensional three-species five-reaction zero-one networks.
-Our purpose is to show the improvement Algorithm 1 makes comparing with purely computing Gröbner bases.
-One can improve the computational timings presented in Table 1 by applying any more advanced tools for computing Gröbner bases rather than the standard Maple package.
-Table 1: Computational time (min: minutes; s: seconds)
+Note that the number of all three-dimensional three-species five-reaction zero-one networks is $3819816$ [16].
+If we do not apply Algorithm 1, then we need about $60$ hours to check the reduced Gröbner bases of these networks.
+Our main purpose is to demonstrate the improvements achieved by Algorithm 1 compared with the pure computation of Gröbner bases.
+The computational time of Step 7 presented in Table 1 can be improved by applying any more advanced tools to calculate the Gröbner bases rather than the standard Maple package.
+Recall that the original version of Algorithm 1, as presented in [16], requires approximately one hour to complete the entire process (see [16, Table 1]).
+Here, the improved version has saved about half of the total computational time by operating on stoichiometric matrices rather than networks in the first three steps.
+Another observation is that most inconsistent networks are filtered out when the vacuous ideals are verified at Step 4.
+Table 1: Computational results for $(3,5,3)$-networks (min: minutes; s: seconds)
+STEP NUMBER OF Matrices TIME
+Step 1 43520 1 min
+Step 2 14444 1 min
+Step 3 2548 1 min
 STEP NUMBER OF NETWORKS TIME
-Step 1 3819816 1 min
-Step 2 484477 30 min
-Step 3 76752 145 s
-Step 4 40522 40 min
-Step 5 32668 100 s
-Step 6 32394 30 min
+Step 4 91802 1 min
+Step 5 48703 30 s
+Step 6 38249 2 min
+Step 7 38185 22 min
 Notes: (i) The column “STEP” lists the steps in the procedure shown in Algorithm 1.
@@
 We run the procedure by a 3.60 GHz Inter Core i9 processor (64GB total memory) under Windows 10.
-Acknowledgments We thank Professor Balázs Boros for his suggestion on how to efficiently generate small networks in the Joint Annual Meeting KSMB-SMB 2024.
+In addition, we apply Algorithm 1 to two key classes of zero-one networks: three-dimensional (with $3$ species and $6$ reactions) and four-dimensional (with $4$ species and $5$ reactions).
+Given that both classes include a large number of networks (the numbers of matrices at Step 1 are $276382$ and $7818144$ respectively), we restrict our computations to quadratic networks only.
+Notably, Algorithm 1 can be easily adjusted to compute quadratic zero-one networks (in Step 1, we only enumerate the stoichiometric vectors for the quadratic reactions and in Step 4, we only consider the reactant matrices $\mathcal{Y}$} for quadratic networks).
+We record the timings for the completion of the steps of Algorithm 1 in Table 2 and Table 3, respectively.
+Another interesting finding is that there exists no networks after carrying out Step 6 for these two classes of quadratic networks.
+That means these networks have no positive steady states for a generic choice of rate constants.
+We show an example of three species six-reaction three-dimensional zero-one quadratic network remaining in Step 5, see Example 7.
+Table 2: Computational results for $(3,6,3)$-quadratic networks (min: minutes; s: seconds)
+STEP NUMBER OF Matrices TIME
+Step 1 10693 1 min
+Step 2 5628 1 min
+Step 3 960 1 min
+STEP NUMBER OF NETWORKS TIME
+Step 4 1897 1 min
+Step 5 1565 0.2 s
+Step 6 0 0.14 s
+Step 7 NULL NULL
+Table 3: Computational results for $(4,5,4)$-quadratic networks (min: minutes; s: seconds)
+STEP NUMBER OF Matrices TIME
+Step 1 1440168 3 min
+Step 2 112008 1 min
+Step 3 4822 5 min
+STEP NUMBER OF NETWORKS TIME
+Step 4 21440 1 min
+Step 5 5032 1.9 s
+Step 6 0 2.9 s
+Step 7 NULL NULL
+Example 7.
+Consider the following $3$-dimensional quadratic zero-one network with $3$ species and $6$ reactions:
+$X_{2}+X_{3}\xrightarrow{\kappa_{1}}X_{1}+X_{3},$ $X_{1}+X_{3}\xrightarrow{\kappa_{2}}X_{2}+X_{3},$ $X_{1}+X_{3}\xrightarrow{\kappa_{3}}X_{1}+X_{2},$
+$X_{1}+X_{2}\xrightarrow{\kappa_{4}}X_{1},$ $X_{1}+X_{2}\xrightarrow{\kappa_{5}}X_{2}+X_{3},$ $X_{1}+X_{2}\xrightarrow{\kappa_{6}}X_{1}+X_{2}+X_{3}.$
+The steady-state system $f$ is
+$f_{1}$ $=\kappa_{1}x_{2}x_{3}-\kappa_{2}x_{1}x_{3}-\kappa_{5}x_{1}x_{2},$
+$f_{2}$ $=-\kappa_{1}x_{2}x_{3}+\kappa_{2}x_{1}x_{3}+\kappa_{3}x_{1}x_{3}-\kappa_{4}x_{1}x_{2},$
+$f_{3}$ $=-\kappa_{3}x_{1}x_{3}+\kappa_{5}x_{1}x_{2}+\kappa_{6}x_{1}x_{2}.$
+From $f_{1}=f_{2}=f_{3}=0$, we can get
+$f_{1}+f_{2}+f_{3}=(-\kappa_{4}+\kappa_{6})x_{1}x_{2}=0.$
+Hence, if $\kappa_{4}=\kappa_{6}$, the network admits infinitely many positive steady states that satisfy
+$x_{1}=\frac{\kappa_{1}\kappa_{3}}{\kappa_{3}\kappa_{5}+\kappa_{2}\kappa_{5}+\kappa_{2}\kappa_{6}}x_{3},\;x_{2}=\frac{\kappa_{3}}{\kappa_{5}+\kappa_{6}}x_{3}.$
+However, the reduced Gröbner basis of the steady-state ideal $\langle f\rangle$ in ${\mathbb{Q}}(\kappa)[x]$ is
+$\{x_{2}x_{3},\ x_{1}x_{3},\ x_{1}x_{2}\}.$
+That means the network indeed has no positive steady states for any generic choice of $\kappa\in{\mathbb{R}}_{>0}^{6}$.
+So, this network is consistent and has a non-vacuous steady-state ideal—hence its passage through Steps 2 and 4 of Algorithm 1—yet it is still dynamically trivial.
+Acknowledgements.
+We thank Professor Balázs Boros for his suggestion on how to efficiently generate small networks in the Joint Annual Meeting KSMB-SMB 2024.
 We thank Professor Murad Banaji for generating and providing all three-dimensional three-species five-reaction zero-one networks by his own efficient procedure as a reference for us.
+We would like to thank the anonymous reviewers of ISSAC 25 for their insightful and valuable comments, which have greatly contributed to the improvement of this work.